### v1

# Co-evolution of Nuclear Star Clusters and Massive Black Holes: Extreme Mass-Ratio Inspirals

Abstract
We explore extreme mass-ratio inspirals (EMRIs) in the co-evolution of massive black holes (MBHs) and nuclear star clusters (NSCs), which host diverse stellar populations across a wide range of masses. The dynamics are simulated self-consistently with GNC, which we have updated to incorporate gravitational wave orbital decay, the loss cone of a spinning MBH, and stellar evolution. Over $12$ Gyr, we investigate the evolution of the NSC with a mass-growing MBH, as well as the EMRIs of stellar black holes, neutron stars, white dwarfs, brown dwarfs (BDs), and low-mass main-sequence stars (MSs), along with tidal disruption events (TDEs) involving MSs, BDs, and post-MSs. The mass growth of the MBH contributed by TDEs is typically $\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. Between $40\%$ and $70\%$ of the stellar mass is lost during stellar evolution, which dominates the mass growth of the MBH if a significant fraction of the lost mass is accreted. The evolution of EMRI rates is generally affected by the cluster’s size expansion or contraction, stellar population evolution, MBH mass growth, and the stellar initial mass function. The EMRI rates for compact objects peak at early epochs ($\lesssim 1$ Gyr) and then gradually decline over cosmic time. LISA-band ($0.1$ mHz) EMRIs involving compact objects around Milky-Way-like MBHs tend to have high eccentricities, while those around spinning MBHs preferentially occupy low-inclination (prograde) orbits. In contrast, MS- and BD-EMRIs usually have eccentricity and inclination distributions that are distinct from those of compact objects.
Black-hole physics – gravitation – gravitational waves – Galaxy: center – Galaxy: nucleus – relativistic processes – stars: kinematics and dynamics
[0]
1 Introduction
An extreme-mass ratio inspiral (EMRI) occurs when a stellar object is gravitationally captured by a central massive black hole (MBH) and continuously spirals inward through the emission of gravitational waves (GWs) (e.g. Hils & Bender, 1995). Due to the strong tidal force of the MBH, such objects are expected to be compact, e.g., stellar-mass black holes (SBHs), neutron stars (NSs), and white dwarfs (WDs) (e.g., Hopman & Alexander, 2005, 2006a, 2006b; Bar-Or & Alexander, 2016; Amaro-Seoane, 2018; Broggi et al., 2022; Qunbar & Stone, 2024), or low-mass main-sequence stars (MSs) and brown dwarfs (BDs, with masses in the range of $0.01M_{\odot}$-$0.1M_{\odot}$(Freitag, 2003; Amaro-Seoane, 2019). For MBHs with masses of $10^{4}M_{\odot}$-$10^{7}M_{\odot}$, an EMRI is a slow process that usually involves hundreds of thousands of inspirals emitting GWs in the frequency range of $10^{-4}$ Hz to $0.1$ Hz. Thus, EMRIs are one of the main targets for future space-based GW observatories, such as the Laser Interferometer Space Antenna (LISA) (Amaro-Seoane et al., 2017; Barausse et al., 2020; Amaro-Seoane et al., 2023), TianQin (e.g., Luo et al., 2016), or TaiJi (Ruan et al., 2020).
The extended duration of the inspiral phase is a crucial characteristic of EMRIs. Because the timescale of gravitational wave emission scales strongly with the semi-major axis ($T\propto a^{4}$), EMRIs spend the majority of their lifetime in the early stages of the inspiral. This leads to the concept of Early EMRIs (E-EMRIs), which can spend hundreds of thousands of years in the detection band before the final merger (Amaro Seoane et al., 2024). Depending on how rapidly their frequency evolves during an observation period, E-EMRIs can be classified as monochromatic (negligible frequency evolution), oligochromatic (slow frequency evolution), or polychromatic (rapid evolution near merger) (Amaro Seoane et al., 2024). While the merger event rate might be low, the long residence time implies that a significant steady-state population of E-EMRIs may be present in the detector band at any given time.
EMRIs form in nuclear star clusters (NSCs) around an MBH when relaxation processes (particularly two-body relaxation) scatter their orbits into regimes where the dynamics are dominated by GW radiation (e.g., Hopman & Alexander, 2005), or when stellar objects are trapped within an accretion disk (e.g., Miralda-Escudé & Kollmeier, 2005; Kocsis et al., 2011; Pan et al., 2021; Pan & Yang, 2021). The GW signals from EMRIs encode detailed information about the MBH mass, spin, and spacetime structure (e.g., Sopuerta & Yunes, 2011). Thus, EMRIs are expected to be ideal tools for probing stellar dynamics in NSCs, black hole physics, and theories of gravity in strong fields (e.g., Ryan, 1995; Gair et al., 2013; Amaro-Seoane, 2018; Cárdenas-Avendaño & Sopuerta, 2024).
Furthermore, the extremely large mass ratios involved when considering the inspiral of brown dwarfs (BDs) around an MBH—termed X-MRIs—lead to even longer residence times, potentially millions of years, suggesting a non-negligible population of these sources might also be detectable (Amaro-Seoane, 2019).
Despite extensive previous studies (e.g., Hils & Bender, 1995; Sigurdsson & Rees, 1997; Alexander & Hopman, 2003; Hopman & Alexander, 2005, 2006a, 2006b; Preto & Amaro-Seoane, 2010; Amaro-Seoane & Preto, 2011; Merritt et al., 2011; Bar-Or & Alexander, 2016; Babak et al., 2017; Raveh & Perets, 2021; Broggi et al., 2022; Qunbar & Stone, 2024; Mancieri et al., 2025), significant uncertainties remain regarding the dynamics, evolution, and predicted properties of EMRIs. Most of the difficulties come from the challenges of numerically simulating the complex stellar dynamics of an NSC around an MBH. For example, an NSC is a dense and compact cluster typically consisting of tens of millions of stellar objects with a broad spectrum of masses. The orbital evolution is mainly driven by two-body stochastic relaxation processes, which are tightly coupled among components of different masses (e.g., Freitag & Benz, 2001, 2002; Freitag et al., 2006; Zhang & Amaro Seoane, 2024).
For accurate simulations of EMRIs, it is also important to distinguish them from ”plunge events,” where objects directly enter the loss cone region, producing only a single burst of GW emission (Hils & Bender, 1995). Even near the loss cone, it is possible for gravitational encounters with background stellar populations to scatter the objects back to less eccentric orbits, interrupting the inspiral process.
The dynamics are further complicated by the spin of the MBH (Amaro-Seoane et al., 2013), resonant relaxation (Rauch & Tremaine, 1996; Bar-Or & Alexander, 2016), relativistic orbital precessions (leading to the so-called ”Schwarzschild barrier”) (e.g., Merritt et al., 2011), and binary effects (e.g., Chen & Han, 2018; Raveh & Perets, 2021). Many environmental effects can contaminate EMRI signals, such as the dynamical perturbations of a secondary MBH (Yunes et al., 2011), background stellar perturbations (Amaro-Seoane et al., 2012), disks (Zwick et al., 2025; Kocsis et al., 2011), and dark matter (e.g., Dai et al., 2024).
As a consequence of these dynamical complexities, many previous studies have adopted various simplifications in the dynamics of EMRIs. For example, the dynamics of EMRIs are often estimated according to the steady-state solution of the system (e.g., Hopman & Alexander, 2005, 2006b; Bar-Or & Alexander, 2016). Many previous studies assume a non-evolving MBH with a fixed mass, ignoring the evolution of stellar populations and the stellar potential across the cluster. The differences in the estimated EMRI event rates are large; for example, SBH-EMRI event rates range from $10^{-9}$ yr-1 (Hopman & Alexander, 2005) to $10^{-6}$ yr-1 (Freitag, 2003).
Another important source of uncertainty is that current observational constraints on the properties of NSCs remain poor, including the mass and density distributions of stellar objects (e.g., Neumayer et al., 2020), except for a few NSCs in nearby galaxies and our own Galactic Center. Thus, many previous studies adopt an NSC structure where the influence radius of the MBH follows the $M-\sigma$ relation (Kormendy & Ho, 2013), under the assumption of an isothermal density distribution for galactic nuclei. However, this assumption becomes inaccurate in several scenarios, such as when the density profile deviates from isothermal, the MBH mass is significantly smaller than the NSC mass (Alexander, 2017; Zhang & Amaro Seoane, 2025), or the cluster exhibits rapid size expansion over cosmic time (Zhang & Amaro Seoane, 2025).
For an accurate study of the dynamics of an NSC and the formation of EMRIs, it is important to simulate the co-evolution of the MBH and the NSC over cosmic time, i.e., the MBH grows its mass by accreting surrounding gaseous material, while the stellar structure of the NSC changes in response to the mass growth of the MBH (e.g., Young, 1980; Freitag & Benz, 2001; Zhang & Amaro Seoane, 2025). Due to stellar evolution, the stellar populations of an NSC should also evolve with time (e.g., Panamarev et al., 2019). The mass loss due to stellar evolution should contribute significantly to the mass growth of the central MBH (Murphy et al., 1991; Freitag & Benz, 2002).
Here we aim to improve the study of EMRIs in NSCs based on our previously well-tested Monte Carlo method, GNC (Zhang & Amaro Seoane, 2024, 2025, Hereafter Paper I and Paper II, respectively). GNC can simulate the two-body relaxation processes by solving the Fokker-Planck equations in the space of orbital energy and angular momentum for various types of stellar objects with a spectrum of masses. GNC integrates the stellar potential and adiabatic invariant theory, allowing for the simulation of the self-consistent dynamics of an NSC with a mass-growing MBH (Paper II). For a more comprehensive study of EMRIs, in this work we additionally include recipes for GW orbital dissipation, the loss cone of a spinning MBH, and stellar evolution processes into GNC. This framework enables the simultaneous modeling of EMRIs, the evolution of NSCs, and the mass growth of MBHs within a unified physical description, a combination not achieved in many previous studies.
As the first application of GNC to EMRI studies, we adopt an idealized scenario of an isolated, gas-free NSC in a galactic nucleus. We assume that all stars form simultaneously in a single burst with no subsequent star formation. While our framework can, in principle, accommodate continuous or multiple bursts of star formation, doing so introduces additional complexities and requires specific assumptions about both the star formation history and the location of star formation within the cluster. We do not consider resonant relaxation here, as its effects on EMRI event rates remain controversial (e.g., Merritt et al., 2011; Bar-Or & Alexander, 2016; Hopman & Alexander, 2006b). We therefore defer investigations of these more complex scenarios, including gaseous environments, the history of star formation, and other physical processes, to future studies.
The paper is organized as follows. In Section 2, we describe how the additional recipes are included in GNC. In Section 3, we perform simulations and investigate the evolution of the NSC, the mass growth of the MBH, and the expected properties of EMRIs in the LISA band. Discussions and conclusions are given in Section 4 and 5, respectively.
2 The method
[…]
We adopt Dehnen’s model (Dehnen, 1993) for the initial density profile of the cluster:
$\rho(r)=\frac{3-\gamma}{4\pi}\frac{M_{\rm cl}r_{a}}{r^{\gamma}(r+r_{a})^{4-\gamma}}$ (1)
where $M_{\rm cl}$ is the total mass of the cluster, $\gamma$ is the density profile in the inner regions of the cluster, and $r_{a}$ marks the characteristic radius where the density profile changes. Then, approximately $10^{5}$ Monte Carlo sample particles are generated in the space of dimensionless energy $x=E/\sigma_{0}^{2}$ and dimensionless angular momentum $j=J/J_{c}(x)$ for given masses and stellar types, where
  • $E=\phi(r)-\frac{1}{2}v^{2}$ is the specific orbital energy, where $\phi(r)$ is the gravitational potential (including the stellar and MBH potential), and $r$ and $v$ are the radial distance to the cluster’s center and the velocity, respectively;
  • $J=rv_{t}$ is the specific orbital angular momentum, where $v_{t}$ is the tangential velocity;
  • $\sigma_{0}$ is the user-defined characteristic velocity. We set $\sigma_{0}=\sqrt{m_{0}G/r_{0}}$, where $G$ is the gravitational constant. $m_{0}$ is the characteristic mass, set to $m_{0}=M_{\bullet}$ if the MBH mass $M_{\bullet}$ is fixed, or $m_{0}=10^{7}M_{\odot}$ if not. $r_{0}$ is the characteristic distance given by $r_{0}=3.1~{\rm pc}(m_{0}/4\times 10^{6}M_{\odot})^{-0.55}$;
  • $J_{c}(x)$ is the maximum angular momentum for a given orbital energy $x$.
The potential $\phi(r)$ as a function of radius $r$ is obtained self-consistently by solving the Poisson equation through iterative methods (see Section 2.2.2 of Paper II). According to the conservation of radial action, the particles’ energies are adjusted in response to the slowly varying stellar and MBH potential (see Section 2.3 of Paper II).
A gas reservoir formed from the mass of TDEs (Zhang & Amaro Seoane, 2025) or from mass loss during stellar evolution (see Section 2.4 and the beginning of Section 3.2) is assumed to be accreted by the MBH. The consumption rate of this gas reservoir is limited by the Eddington mass accretion rate:
$\dot{M}_{\rm edd}=2.22~M_{\odot}~{\rm yr}^{-1}\times\frac{0.1}{\epsilon}\frac{M_{\bullet}}{10^{8}M_{\odot}},$ (2)
where $\epsilon=0.1$ is the radiative efficiency. The mass growth of the MBH is then given by $\dot{M}_{\bullet}=\dot{M}_{\rm edd}(1-\epsilon)$.
The simulation is performed in iterations with adaptive timesteps $\Delta t$, which is the minimum of the two-body relaxation timescale and the accretion timescale when modeling the mass growth of the MBH (see Equation 40 of Paper II).
The diffusion coefficients for the two-body relaxation process—i.e., the drift ($D_{E}^{\rm NR}$ and $D_{J}^{\rm NR}$), diffusion ($D_{EE}^{\rm NR}$ and $D_{JJ}^{\rm NR}$), and cross ($D_{EJ}^{\rm NR}$) terms for energy and angular momentum, which include the effects of the stellar potential—are pre-calculated before each iteration. Within each simulation iteration of timestep $\Delta t$, each individual particle iteratively evolves its energy and angular momentum with smaller adaptive timesteps $\delta t$ (Equation B4 of Paper I):
$\delta E$ $=D_{E}\delta t+y_{1}\sqrt{D_{EE}\delta t},$ (3)
$\delta J$ $=D_{J}\delta t+y_{2}\sqrt{D_{JJ}\delta t}.$
where $D_{E}$ ($D_{J}$) is the sum of the drift terms including $D_{E}^{\rm NR}$ ($D_{J}^{\rm NR}$), and $D_{EE}=D_{EE}^{\rm NR}$ ($D_{JJ}=D_{JJ}^{\rm NR}$) is the diffusion term for energy (angular momentum). $y_{1}$ and $y_{2}$ are two random numbers drawn from a standard normal distribution with correlation $D_{EJ}^{\rm NR}/\sqrt{D_{EE}^{\rm NR}D_{JJ}^{\rm NR}}$.
A particle is removed if its energy is so high that it moves too close to the MBH (typically $x>10^{5}-10^{6}$) or if its energy is too low (typically $x\lesssim 10^{-3}$) such that it can easily escape from the system. If the loss cone is included, particles with $j<j_{\rm lc}$ are also removed, where $j_{\rm lc}$ is the size of the dimensionless angular momentum of the loss cone. After all particles are evolved, the self-consistent potential, density, and phase-space distribution of $x$ and $j$ are updated for the next iteration.
[…]
2.1 Including orbital decay by gravitational wave radiation
Gravitational wave emission from a stellar object around an MBH is usually considered for orbits bound to the MBH. However, since we have included the stellar potential, the orbit of a stellar object may be bound to the NSC but unbound to the MBH. If the orbit of such an object approaches the loss-cone regime, it encounters the MBH on a hyperbolic orbit. In this case, the resulting gravitational wave emission differs slightly from that of a bound orbit.
When the stellar potential is included, Keplerian orbital elements are no longer accurate, as they are defined in a two-body system. Nevertheless, since gravitational wave dissipation is most important near the pericenter, the GW emission can still be accurately estimated using the instantaneous Keplerian semi-major axis $a_{K}$ and eccentricity $e_{K}$ at the pericenter $r_{p}(E,J)$, where $r_{p}(E,J)$ is the pericenter of the orbit with energy $E$ and angular momentum $J$ (obtained by solving Equation 5 of Paper II). Since $a_{K}(1-e_{K})=r_{p}(E,J)$, we can derive the instantaneous values of $e_{K}$ and $a_{K}$:
$e_{K}$ $=\left|\frac{J^{2}}{M_{\bullet}Gr_{p}(E,J)}-1\right|$ (4)
$a_{K}$ $=\frac{r_{p}(E,J)}{1-e_{K}}.$
For a particle bound (unbound) to the MBH, near the pericenter we instantaneously have $e_{K}<1$ and $a_{K}>0$ ($e_{K}>1$ and $a_{K}<0$).
The GW orbital decay can be added to the drift terms for orbital energy and angular momentum in Equation 3. To model EMRI events self-consistently, we calculate the GW radiation for bound or unbound orbits separately. The details are in the following two sections.
2.1.1 For orbits bound to the MBH
Let $D^{\rm GW}_{E}$ ($D^{\rm GW}_{J}$) be the drift of energy (angular momentum) due to GW radiation. For stellar orbits bound to the MBH, the orbit-averaged decay of energy and angular momentum per orbit, $\delta E^{\rm GW}$ and $\delta J^{\rm GW}$, are given by (Peters & Mathews, 1963; Peters, 1964):
$\delta E^{\rm GW}=D^{GW}_{E}P$ (5)
$=$ $-\frac{64\pi}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}}\frac{1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}}{(1+e_{K})^{7/2}},$
$\delta J^{\rm GW}=D^{GW}_{J}P$
$=$ $-\frac{64\pi}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}}\frac{1+\frac{7}{8}e^{2}_{K}}{(1+e_{K})^{2}},$
respectively, where $c$ is the speed of light, $m$ is the mass of the particle, and $P$ is the orbital period.
In reality, the GW dissipation of both $E$ and $J$ varies along the orbit, with the maximum and minimum occurring near the pericenter and apocenter, respectively. For moderately eccentric orbits (e.g., $e\lesssim 0.9$), the variation is relatively mild, and thus the change in energy and angular momentum within a given time $\delta t$ can be approximated by $D^{\rm GW}_{E}\delta t$ and $D^{\rm GW}_{J}\delta t$, respectively.
However, for highly eccentric orbits (e.g., $e\gtrsim 0.9$), GW dissipation is strongly concentrated near the pericenter, with negligible dissipation elsewhere along the orbit. In these cases, applying a smooth damping term would lead to a significant overestimation of the dissipation before the pericenter passage and an underestimation after it. This is particularly severe for particles weakly bound to the MBH but with extremely high orbital eccentricity, as before pericenter passage, the particle’s dynamics may actually be dominated by relaxation rather than GW radiation.
Therefore, for a particle with high eccentricity ($e>0.9$), in each orbital revolution, we apply the full orbital changes $\delta E^{\rm GW}=D^{\rm GW}_{E}P$ and $\delta J^{\rm GW}=D^{\rm GW}_{J}P$ only if it has passed the pericenter.
2.1.2 For orbits unbound to the MBH
For stellar objects unbound to the MBH but bound to the cluster, after each pericenter passage, the decay of orbital energy and angular momentum are given by (Turner, 1977; Hansen, 1972):
$\begin{gathered}\delta E^{\rm GW}=D^{GW}_{E}P=-\frac{64}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}(1+e_{K})^{7/2}}\\ \times\left[\left(\pi-\theta_{0}\right)\left(1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}\right)\right.\\ +\left.\frac{(e^{2}_{K}-1)^{1/2}}{144}\left(301+\frac{673}{2}e^{2}_{K}\right)\right],\\ \delta J^{\rm GW}=D^{GW}_{J}P=-\frac{64}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}(1+e_{K})^{2}}\\ \times\left[(\pi-\theta_{0})\left(1+\frac{7}{8}e^{2}_{K}\right)+\frac{e_{K}}{8}\sin\theta_{0}(13+e^{2}_{K})\right],\end{gathered}$ (6)
respectively, where $\cos\theta_{0}=1/e_{K}$. Similarly to Section 2.1.1, here we damp the energy and angular momentum only if the particle has passed the pericenter of its orbit.
2.2 Identifying EMRIs in the simulation
In our simulations, a particle is safely identified as an EMRI if its GW orbital dissipation timescale is much shorter than the two-body relaxation timescale. Specifically, this condition is satisfied when
$T_{\rm GW}(E,J)=\frac{E}{|D^{\rm GW}_{E}(E,J)|}$ (7)
$\leq$ $CT_{\rm rlx}(E,J)=C{\rm min}\left(\frac{J^{2}}{D_{JJ}^{\rm NR}(E,J)},\frac{E^{2}}{D_{EE}^{\rm NR}(E,J)}\right)$
where $T_{\rm GW}(E,J)$ is the GW orbital decay time and $T_{\rm rlx}(E,J)$ is the two-body relaxation time for a given orbital energy $E$ and angular momentum $J$. $D_{E}^{\rm GW}$ is the energy damping rate given by either Equation 5 for bound orbits or Equation 6 for unbound orbits. $D_{JJ}^{\rm NR}$ and $D_{EE}^{\rm NR}$ are the diffusion coefficients from Equations 21 and 23 in Paper II, respectively. $C$ is a threshold parameter specifying the ratio of the two-body timescale to the GW decay timescale.
We set $C=10^{-3}$, which is sufficiently small so that the obtained EMRI event rate converges. This conservative threshold prevents the false identification of EMRI events. Thresholds with $C\gtrsim 0.1$ are unreliable, as particles moving in the regime $0.001T_{\rm rlx}\lesssim T_{\rm GW}\lesssim 0.1T_{\rm rlx}$ can possibly be scattered back to $T_{\rm GW}>0.1T_{\rm rlx}$ by the stochastic two-body relaxation process (see Figure 2 for trajectories of particles in regimes where $10^{-3}T_{\rm rlx}<T_{\rm GW}<0.1T_{\rm rlx}$).
Particles that fall directly into the loss cone from regimes not satisfying Equation 7 are identified as plunge events. For more details on the simulated trajectories of EMRIs and plunge events, see Section 3.1.1.
2.3 Loss cones around a spinning MBH
The event rates of EMRIs depend critically on the size of the loss cone. The loss cone is the regime in the space of orbital pericenter $r_{p}<r_{\rm p,lc}$ (or angular momentum $j<j_{\rm lc}$) where a particle can be destroyed or swallowed by the MBH. Generally, $r_{\rm p,lc}={\rm max}(r_{\rm ISO},r_{\rm td})$ is the maximum of the pericenter of the Innermost Stable Orbit (ISO), $r_{\rm ISO}$, and the tidal radius, $r_{\rm td}$. In the space of angular momentum, it is $j_{\rm lc}={\rm max}(j_{\rm ISO},j_{\rm td})$, where $j_{\rm ISO}$ and $j_{\rm td}$ are the angular momenta corresponding to $r_{\rm ISO}$ and $r_{\rm td}$, respectively. For compact objects, $r_{\rm p,lc}=r_{\rm ISO}$ or $j_{\rm lc}=j_{\rm ISO}$. However, for gaseous stars, such as BDs and MSs, the additional limitation by the tidal radius must be included.
If the MBH is spinning, the size of the ISO, $r_{\rm ISO}$, varies with the orbital inclination $i$ of the particle relative to the equatorial plane of the MBH. Following a method similar to that in Sopuerta & Yunes (2011) (see also Amaro-Seoane et al. (2013)), we derive the pericenter of the ISO, $r_{\rm ISO}$, for a particle in an inclined orbit around a spinning MBH. Generally, $r_{\rm ISO}=\mathcal{W}r_{g}$, where $\mathcal{W}$ depends on the dimensionless spin parameter $a$ and inclination $i$. For a Schwarzschild MBH ($a=0$), $\mathcal{W}=8$. If $a=1$, $\mathcal{W}$ varies from $2.12$ (prograde orbits, $i=0^{\circ}$) to $11.65$ (retrograde orbits, $i=180^{\circ}$). For more details, see Appendix A and B.
In reality, the orbital orientation should evolve, e.g., due to vector resonant relaxation (Rauch & Tremaine, 1996). However, the spatial orientation of the orbit is likely to remain randomized during stochastic relaxation processes. Thus, for simplicity, we assign an initially random orbital inclination to each particle, which remains fixed over time.
2.4 Stellar evolution
Figure 1: The mass-radius relation for BDs ($m_{\star}=0.01M_{\odot}\sim 0.1M_{\odot}$) and low-mass MSs ($0.1M_{\odot}\sim 1M_{\odot}$). Black lines are results from MOBSE, while colored lines are from Baraffe et al. (2003). The green dashed lines show the stellar radius at which the tidal radius of the star, $r_{\rm td}$, equals the pericenter of the ISO ($r_{\rm ISO}$). For prograde (retrograde) orbits around a maximally spinning MBH, $r_{\rm ISO}=2.12r_{g}$ ($r_{\rm ISO}=11.65r_{g}$). For a Schwarzschild MBH, $r_{\rm ISO}=8r_{g}$.
Real NSCs consist of stellar objects of multiple types, including MSs, post-main-sequence stars (post-MSs), and compact objects. To accurately simulate an NSC’s evolution, we incorporate stellar evolution using MOBSE (Giacobbo et al., 2018; Giacobbo & Mapelli, 2018), which is an upgraded version of BSE (Hurley et al., 2000, 2002). The stellar evolution in MOBSE is the same as that in BSE but includes up-to-date equations for metal-dependent stellar winds, new prescriptions for core-collapse supernovae, and a dependence of stellar winds on the Eddington factor. Moreover, MOBSE extends the mass range of stars from $100M_{\odot}$ to $150M_{\odot}$. For simplicity, in this study we ignore any natal kicks for NSs and SBHs.
We assume a single-starburst formation scenario, where all stellar objects begin as zero-age MSs. The evolution of each MS is pre-computed using MOBSE over a $12$ Gyr period. During the simulation, the stellar type and radius of each particle are updated according to its evolutionary table.
The stellar evolution results in a substantial population of post-MSs, including objects in the phases of the Hertzsprung gap, first giant branch, core helium burning, and first and second asymptotic giant branches. For simplicity, we collectively classify them as ”Post-MSs” and do not distinguish individual subtypes in our simulation. The evolution also generates various naked helium stars. However, as their total number is negligible, we include them in the simulation without further analysis. Finally, although MOBSE resolves multiple types of WDs (Helium, Carbon/Oxygen, and Oxygen/Neon WDs), we group all of them together as ”WDs,” as we are not currently interested in investigating individual WD subtypes.
To accurately track the evolution, the simulation timestep must resolve key evolutionary transitions. Specifically, if $\delta t$ is the update timestep for each particle (see Appendix B in Paper I), then we impose an additional constraint $\delta t\rightarrow\min(\delta t,t_{\rm nk}-t)$, where $t$ is the current time of the particle and $t_{\rm nk}$ is the next key evolutionary transition time from the pre-calculated table.
2.5 The mass-radius relation and IMF of brown dwarfs
We note that both MOBSE and BSE use an oversimplified mass-radius relation for BDs, i.e., $r_{\star}\propto m_{\star}^{-1/3}$ for $m_{\star}<0.1M_{\odot}$ (Hurley et al., 2000). The stellar radii of BDs are overestimated by a factor of about 232-3 compared to those in Baraffe et al. (2003) (see Figure 1). As the rate of EMRIs is sensitive to the stellar radius of BDs, such an overestimation of the stellar radius increases the probability of their tidal disruption and thus suppresses the formation rate of BD-EMRIs. To improve the modeling of BD-EMRIs, we update each BD’s radius in the simulation according to the age-dependent radius-mass relation for BDs from Baraffe et al. (2003).
For simplicity, in this work we assume that the IMFs of BDs and MSs are continuous around $\sim 0.1M_{\odot}$. In this case, the number ratio of BDs to MSs for all models in this work is $\sim 0.9$. We note, however, that BDs may constitute a distinct population from MSs (Kroupa et al., 2013), resulting in a discontinuous IMF between the two. If we instead adopt such an IMF for BDs, the number ratio of BDs to MSs decreases to $\sim 0.2$. Thus, our simulated event rates for BD-related phenomena, such as EMRIs or tidal disruption events (TDEs) of BDs, would be reduced by a factor of $\sim 4.5$. Nevertheless, since BDs contribute only a small fraction of the total cluster mass ($\sim 5\%$), the majority of the simulation results, including the evolution of NSCs, the mass growth of the MBH, and EMRIs of other objects, are not significantly affected by adopting a continuous or discontinuous IMF for BDs.
3 The dynamics of EMRIs in evolving NSCs
Table 1: Models with discrete mass components
Name $r_{a}^{a}$ $\gamma^{a}$ $r_{\rm eff,i}^{b}$ $r_{\rm eff,f}^{c}$ $r_{\rm h,f}^{c}$ $M_{\bullet}^{d}$ ($M_{\odot}$) Componentse
M2 $2.17$ 11 $3.9$ $6.7$ $3.9$ $4\times 10^{6}$, Fix
MSs ($1M_{\odot}$)
[…]
M5 $2.17$ 11 $3.9$ $6.7$ $3.8$
MSs ($1M_{\odot}$)
[…]
M5_2 $2.91$ $1.5$ $3.9$ $6.6$ $3.8$
M5G82 $1.5$ 11 $2.7$ $4.4$ $2.7$ $10^{4}$, Growth
Figure 2: Examples of evolutionary trajectories of SBHs in model M2 (see Table 1) simulated by GNC, plotted in the space of energy, angular momentum, and pericenter distance. Thin lines ending with filled red circles mark plunge events, while those ending with filled red squares mark EMRI events. Top left panel: The energy and angular momentum evolution for a maximally spinning MBH. $x=E/\sigma_{0}^{2}$ and $j=J/J_{c}(x)$ are the dimensionless energy and angular momentum (see text in Section 2 for their definitions). The loss cone size depends on the amplitude ($a$) and inclination angle (ii) of the MBH’s spin (defined in Equation A4). Top right panel: Similar to the left panel but for the evolution in the space of energy and pericenter distance. $r_{0}=3.1$ pc is a characteristic distance. Bottom panels are similar to the top panels but for the case of a Schwarzschild MBH ($a=0$). The cyan filled triangle, star, and square in each panel mark the intersection between $T_{\rm GW}=0.1T_{\rm rlx}$ and the loss cone for $a=1$, $a=0$, and a=1a=-1 (equivalent to $a=0$ with $i=180^{\circ}$), respectively.
Figure 3: Critical boundaries in the space of dimensionless energy $x$ and angular momentum $j$ for different stellar objects in Model M5. In all panels, black solid lines correspond to $T_{\rm GW}=0.1T_{\rm rlx}$, where $T_{\rm rlx}$ and $T_{\rm GW}$ are the relaxation and GW timescales, respectively (defined in Equation 7). Red, blue, and green dashed lines indicate the pericenter of the ISO: $r_{p}=r_{\rm ISO}=2.12r_{g}$ (for $a=1$ and $i=0^{\circ}$), $r_{\rm ISO}=8r_{g}$ (for $a=0$), and $r_{\rm ISO}=11.6r_{g}$ (for $a=1$ and $i=180^{\circ}$), respectively. The black dashed line in the right panel represents the tidal radius of a BD ($r_{p}=r_{\rm td}=4r_{g}$) for an MBH mass of $4\times 10^{6}M_{\odot}$. The loss cone size is given by $r_{p,\rm lc}={\rm max}(r_{\rm ISO},r_{\rm td})$, and its intersection with $T_{\rm GW}=0.1T_{\rm rlx}$ defines the critical distance $a_{\rm crit}$ (marked by red, blue, green, or black filled circles in each panel). Horizontal dash-dotted lines, colored the same as the filled circles, show the corresponding $a_{\rm crit}$ given by Equation 8.
Figure 4: Evolution of the EMRI and plunge event rates in five-component models (model M5 or M5_2 in Table 1), with a Schwarzschild ($a=0$) or maximally spinning MBH ($a=1$). In these simulations, the MBH mass is fixed, and stellar evolution is ignored.
Simulations incorporating all the aforementioned recipes can provide a more comprehensive study of the evolution of NSCs, the mass growth of the MBH, and the dynamics of EMRIs. However, the coupling between these recipes substantially complicates the analysis of the simulation results. For a clear analysis of EMRI dynamics, we first adopt simplified models in Section 3.1, assuming that the NSC consists of either two or five mass components while neglecting stellar evolution. These simplified scenarios facilitate the analysis of the dynamics in later sections and comparison with many previous studies that adopted similar assumptions. We extend our analysis later in Section 3.2 to more realistic scenarios, where the NSC consists of a spectrum of mass components and includes stellar evolution.
3.1 NSCs with fixed mass components and no stellar evolution
In this section, we ignore stellar evolution and focus on NSC models composed of either two or five components. The initial conditions of the models are shown in Table 1. All models adopt $M_{\rm cl}=4\times 10^{7}M_{\odot}$, similar to that of our Milky Way’s NSC. In the two-component model (M2), the NSC consists of $1M_{\odot}$ MSs and $10M_{\odot}$ SBHs, where SBHs have a number fraction of $0.001$ relative to MSs. In the five-component models (M5, M5_2, or M5G82), we additionally include $0.05M_{\odot}$ BDs with a number fraction of $0.9$ relative to MSs, $0.6M_{\odot}$ WDs with a fraction of $0.2$, and $1.4M_{\odot}$ NSs with a fraction of $0.01$. The fraction of BDs follows the results of $12$ Gyr of continuous star formation with a Kroupa IMF (Kroupa, 2001), as shown in Figure 9 and Table 2 of Paper I. The radius of the BDs is $0.08R_{\odot}$ according to Baraffe et al. (2003) for a $0.05M_{\odot}$ BD with an age of >5>5 Gyr.
We first focus on models M2, M5, and M5_2 (Sections 3.1.13.1.2, and 3.1.3), all of which adopt a fixed MBH mass of $4\times 10^{6}M_{\odot}$, similar to the MBH in our Galactic Center (Gillessen et al., 2009). We then investigate the slightly more complicated model M5G82 (Section 3.1.4), in which the MBH mass grows from $10^{4}M_{\odot}$ by swallowing objects that fall into the loss cone (including EMRIs and plunge events) and by accreting gas released by MS-TDEs. Details of the accretion processes can be found in Section 4.2 of Paper II. For all models, the evolution and rates of TDEs are similar to those in Paper II, and thus we will not discuss them further here.
3.1.1 The formation of EMRIs
Figure 2 shows the evolution of SBHs in the space of dimensionless energy ($x$), angular momentum ($j$), and pericenter distance $r_{p}/r_{0}$ ($r_{0}=3.1$ pc) for the two-component model (M2) around a Schwarzschild or a maximally spinning MBH. It is apparent that below the boundary defined by Equation 7, the evolution is dominated by GW radiation, and the SBHs eventually become EMRIs.
The reduced pericenter distance of ISOs around a maximally spinning MBH —reaching as low as $2.12r_{g}$ compared to the $8r_{g}$ of a Schwarzschild black hole—significantly expands the spatial regime within the cluster where particles can become EMRIs. This enhanced parameter space particularly affects particles in the cluster’s outer regions, enabling them to transition into EMRIs despite their larger orbital semi-major axes. As a consequence, we expect that the rates of EMRIs will be higher around a spinning MBH.
SBHs entering the loss cone from regimes above the boundary become plunge events. It is apparent from the right panels of Figure 2 that they fall into the MBH with a pericenter distance smaller than $r_{\rm ISO}$ (the pericenter of the ISO). Note that a small fraction of plunge events exhibit multiple decreases in orbital energy due to GW emission before falling into the loss cone. However, they are not identified as EMRIs because their dynamics are still dominated by the relaxation process rather than by GWs, i.e., they do not satisfy Equation 7.
3.1.2 The critical radius $a_{\rm crit}$ for different stellar objects
The intersection between $T_{\rm GW}\sim 0.1T_{\rm rlx}$ [^1] Note that here we use $C\sim 0.1$ so that the majority of EMRI events are identified within a relaxation timescale. This is different from the value $C=0.001$ described in Section 2.2, which is the critical value to identify a particle as an EMRI for an instantaneous momentum in the simulation. and the loss cone size, $j=j_{\rm lc}$, defines a critical radius $a_{\rm crit}$ (or energy $E_{\rm crit}$) below (or above) which particles are likely to become EMRIs within their relaxation time (e.g., Hopman & Alexander, 2006a). The critical radius $a_{\rm crit}$ corresponds to the radius of a circular orbit $r_{c}(E_{\rm crit})$ for the critical energy $E_{\rm crit}$. If the stellar potential is ignored, $a_{\rm crit}=GM_{\bullet}/(2E_{\rm crit})$, which reduces to the Keplerian orbital semi-major axis. The larger the value of $a_{\rm crit}$, the higher the expected rates of EMRIs.
The critical radius $a_{\rm crit}$ for different stellar objects can be obtained from our simulation. In Model M5, the positions of $a_{\rm crit}$ for various types of stellar objects are shown in Figure 3. For prograde orbits around a maximally spinning MBH, $a_{\rm crit}\sim 210$ mpc, $43$ mpc, $20$ mpc, and $0.7$ mpc for SBHs, NSs, WDs, and BDs, respectively. These values are $\sim 10$ times larger than those for a Schwarzschild MBH and $\sim 20$ times larger than in the case of retrograde orbits around a maximally spinning MBH. Thus, assuming initially randomized stellar orbit orientations, the EMRI rates around a spinning MBH are expected to be higher than around a Schwarzschild MBH.
Note that for BDs, the loss cone size is limited by their tidal radius. For example, in the case of $a=1$ and $i=0^{\circ}$, we still have $r_{\rm p,lc}=r_{\rm td}=4r_{g}>r_{\rm ISO}=2.1r_{g}$, as shown in the right panel of Figure 3. As a consequence, the enhancement of BD-EMRI event rates around a maximally spinning MBH is not as significant as for other compact objects.
Note that the smaller the MBH mass, the smaller the value of $a_{\rm crit}$ for BDs. When $M_{\bullet}\lesssim 10^{5}M_{\odot}$, BDs become EMRIs only if they can be inside the ISO. In these cases, none of the BDs can become EMRIs.
We can compare $a_{\rm crit}$ from our simulation with results from theoretical analysis, which is approximately given by (Amaro-Seoane, 2019):
$a_{\rm crit}$ $=\epsilon R_{0}\left(\mathscr{W}^{5/2}N_{\bullet}\ln\Lambda\frac{m_{\rm SBH}^{2}}{M_{\bullet}m_{\rm EMRI}}\right)^{\frac{1}{\beta-3}}$ (8)
$\epsilon$ $=\left[16.97C(3-\beta)(1+\beta)^{3/2}\right]^{1/(\beta-3)}$
where
  • $m_{\rm EMRI}$ is the mass of the EMRI object;
  • $m_{\rm SBH}$ is the mass of the SBHs;
  • $\mathscr{W}=r_{p,\rm lc}/(8r_{g})={\rm max}(r_{\rm ISO},r_{\rm td})/(8r_{g})$. If $r_{p,\rm lc}=r_{\rm ISO}$, it depends on the amplitude $a$ and inclination $i$ of the MBH’s spin;
  • $R_{0}$ is the radius within which dynamics are dominated by SBHs, obtained by solving $\rho_{\star}(R_{0})m_{\star}=\rho_{\bullet}(R_{0})m_{\rm SBH}$, where $\rho_{\star}$ and $\rho_{\bullet}$ are the mass density distributions of stars and SBHs;
  • $N_{\bullet}$ is the number of SBHs within $R_{0}$;
  • $C=0.1$; $\Lambda=M_{\bullet}/M_{\odot}$.
  • $\beta$ is the density slope of the objects contributing to EMRIs. According to Paper I, we can approximately adopt $\beta=1.75$ for SBHs and $\beta=1.5$ for other compact objects.
The precise values of both $R_{0}$ and $N_{\bullet}$ in Equation 8 can be obtained from our numerical simulations. For model M5, we find that at $12$ Gyr, $R_{0}=0.26$ pc and $N_{\bullet}=1946$. Adopting these values, we can estimate $a_{\rm crit}$ for different stellar objects from Equation 8. The values for SBHs, NSs, WDs, and BDs are shown as horizontal dash-dotted lines in the panels of Figure 3. We can see that the simulation results are roughly consistent (within about one order of magnitude) with the analytical estimations.
3.1.3 The evolution of the event rates
Figure 4 shows the evolution of EMRI and plunge event rates in models M5 and M5_2, for both a spinning and a non-spinning MBH. The EMRI rates depend on the spin of the MBH and exhibit significant evolution over cosmic time.
For SBHs, the EMRI rates rise and peak around $0.4\sim 1$ Gyr with values of $2.18\sim 6.88\times 10^{-7}$ yr-1, then gradually decline to $\sim(4.7-8.3)\times 10^{-8}$ yr-1 by $12$ Gyr, mainly due to the depletion of SBHs and the size expansion of the cluster. The plunge event rates for SBHs follow a similar evolution but with much higher peak rates of $\sim 2\times 10^{-6}$ yr-1. Such a rise-and-fall behavior of SBH-EMRIs and plunge events—when the mass growth of the MBH is neglected—is consistent with the results in Broggi et al. (2022).
For NSs, WDs, and BDs, their initial rates before $0.1$ Gyr are $(0.1-4)\times 10^{-8}$ yr-1, $(0.1-20)\times 10^{-8}$ yr-1, and $(0.1-10)\times 10^{-8}$ yr-1, respectively, but then decrease gradually to $(0.8-3)\times 10^{-9}$ yr-1, $(0.9-4)\times 10^{-9}$ yr-1, and $(2-8)\times 10^{-10}$ yr-1 by $12$ Gyr, respectively.
These results can be understood from the discussion of $a_{\rm crit}$ in Section 3.1.2. As the $a_{\rm crit}$ of NSs, WDs, and BDs are much smaller than that of SBHs, their rates of EMRIs are orders of magnitude lower. For all compact objects, the EMRI rates are generally $\sim 3-4$ times higher around a maximally spinning MBH than around a non-spinning MBH.
Theoretically, the rates can be estimated using Equation 32 of Amaro-Seoane (2019):
$\dot{\Gamma}_{\rm EMRI}$ $\simeq\frac{3-\beta}{2\lambda}\frac{N_{\rm obj,0}}{T_{0}R_{0}^{\lambda}}a_{\rm crit}^{\lambda}\left[\ln(\Lambda_{\rm crit})-\frac{1}{\lambda}\right]$ (9)
where
  • $\lambda=9/2-\beta-\gamma$, where $\gamma$ is the power-law index of the number density of SBHs;
  • $\Lambda_{\rm crit}=a_{\rm crit}/(8R_{s})$, where $R_{s}$ is the Schwarzschild radius;
  • $N_{\rm obj,0}$ is the number of EMRI objects within $R_{0}$;
  • $T_{0}$ is the two-body relaxation time at $R_{0}$.
The definitions of $a_{\rm crit}$, $R_{0}$, and $\beta$ are the same as in Equation 8. At $12$ Gyr in model M5, the values of $N_{\rm obj,0}$ for SBHs, NSs, WDs, and BDs within $R_{0}$ are $\sim 2\times 10^{3},1.4\times 10^{3},6.4\times 10^{3}$, and $3.5\times 10^{4}$, respectively. In the case of a Schwarzschild MBH, substituting these values into Equation 8, we find that $\Gamma_{\rm EMRI}$ for SBHs, NSs, WDs, and BDs is $\sim 3\times 10^{-8}$ yr-1, $10^{-9}$ yr-1, $4\times 10^{-9}$ yr-1, and $10^{-9}$ yr-1, respectively. These rates are consistent within one order of magnitude with the EMRI rates at $12$ Gyr shown in Figure 4.
For plunge events, all these compact objects have relatively high rates of $10^{-6}\sim 10^{-4}$ yr-1, which depend weakly on the spin of the MBH and the initial density slope of the cluster.
We find that the long-term evolution of EMRI rates is primarily controlled by the mean density of the cluster ($\propto M_{\rm cl}/r_{\rm eff}^{3}$), while an enhanced initial inner density only increases rates temporarily. Figure 4 shows the EMRI evolution in models M5 and M5_2, both of which have the same initial effective size, but the latter has a higher inner density. We can see that both models have almost the same EMRI rates for SBHs, NSs, WDs, and BDs after $\sim 1$ Gyr, although initially the rates in model M5_2 are about one order of magnitude higher.
3.1.4 Including the MBH’s mass growth
Figure 5: Left panel: Evolution in Model M5G82 (see Table 1) of the MBH mass (black solid line), the theoretical mass growth under the Eddington limit (black dashed line), the current mass of the gas reservoir (”Gas reservoir (left)”), and the cumulative contributions from TDEs of stars (”TD (stars)”), SBH-EMRIs (”EMRI (SBHs)”), and plunge events of SBHs and WDs (”Plunge (SBHs)” and ”Plunge (WDs)”). Contributions from plunge/EMRI events of both NSs and BDs are negligible and are thus not shown for clarity. Right panel: Evolution in the same model of the event rates of TDEs of stars or BDs, EMRIs, and plunge events of compact objects.
We also explore the evolution when the MBH mass is growing. The method is similar to that in Section 4.2 of Paper II, but here we additionally include the mass growth of the MBH from swallowing EMRIs. The simulation results for model M5G82 in Table 1 can be found in Figure 5, where the MBH mass grows from $10^{4}M_{\odot}$ to $\sim 3.8\times 10^{6}M_{\odot}$ over $12$ Gyr.
The SBH-EMRI event rates gradually decrease from $4\times 10^{-6}$ yr-1 before $1$ Gyr to $\sim 10^{-7}$ yr-1 by $12$ Gyr. For NSs and WDs, the EMRI rates remain nearly constant at $0.2\sim 2\times 10^{-8}$ yr-1. For BDs, the rates slowly increase from $10^{-10}$ yr-1at $1$ Gyr to $\sim 10^{-9}$ yr-1 by $12$ Gyr. The evolution of EMRI events exhibits different behaviors from those of models with a fixed MBH mass (M5 or M5_2 in Figure 4). Thus, including the mass growth of the MBH is crucial for accurately modeling the cosmological evolution of EMRI events.
We find that the final MBH mass and the evolution of stellar TDEs and SBH plunge events in model M5G82 are quite similar to those in model M2G82 from ( Paper II, see its Figure 10), although the latter considers only two mass components (stars and SBHs).
3.2 NSCs with a mass spectrum, stellar evolution, and stellar mass loss
Table 2: Models with a mass spectrum
Name $M_{\rm cl,i}^{a}$ ($M_{\odot}$) $M_{\rm cl,f}^{a}$ ($M_{\odot}$) $r_{a}$ $r_{\rm eff,i}$ $r_{\rm eff,f}$ Componentsb $f_{\rm ma}^{c}$ $a^{d}$ $M_{\bullet,f}^{e}$ ($M_{\odot}$) $r_{\rm h,f}^{e}$
MKG91F0k $10^{9}$ $5.8\times 10^{8}$ $2.56$ $4.66$ $8.3$
Kroupa IMF
MSs ($0.01M_{\odot}-150M_{\odot}$)
0 11 $3\times 10^{7} 2.2$2.2
MKG91F0.1k $10^{9}$ $5.8\times 10^{8}$ $2.56$ $4.66$ $7.7$ $0.1$ 11 $5.9\times 10^{7}$ $3.7$
MKG81F0k $7\times 10^{7}$ $3.8\times 10^{7}$ $0.9$ $1.6$ $4.9$ 0 11 $4.0\times 10^{6}$ $3.0$
MKG82F0k $7\times 10^{7}$ $4.0\times 10^{7}$ $1.5$ $2.7$ $6.5$ 0 11 $2.0\times 10^{6}$ $2.5$
MKG82F0.06k $7\times 10^{7}$ $3.9\times 10^{7}$ $1.5$ $2.7$ $6.8$ $0.06$ 11 $4.0\times 10^{6}$ $3.9$
MKG82F0.06s $7\times 10^{7}$ $3.9\times 10^{7}$ $1.5$ $2.7$ $7.0$ $0.06$ 0 $\times 10^{6} 3.93.9$
MKG83F0k $7\times 10^{7}$ $4.2\times 10^{7}$ $2.3$ $4.2$ $8.4$ 0 11 $8.6\times 10^{5}$ $2.0$
MKG83F0.1k $7\times 10^{7}$ $4.0\times 10^{7}$ $2.3$ $4.2$ $9.0$ $0.1$ 11 $4.2\times 10^{6}$ $5.1$
MKG83F1k $7\times 10^{7}$ $3.7\times 10^{7}$ $2.3$ $4.2$ $5.2$ 11 11 $2.6\times 10^{7}$ $4.5$
MKG71F0k $2\times 10^{6}$ $1.0\times 10^{6}$ $0.39$ $0.7$ $8.7$ 0 11 $6.3\times 10^{4}$ $4.3$
MKG71F0.1k $2\times 10^{6}$ $0.9\times 10^{6}$ $0.39$ $0.7$ $8.6$ $0.1$ 11 $1.4\times 10^{5}$ $7.1$
MTG82F0.02k $1.7\times 10^{8}$ $4.1\times 10^{7}$ $1.5$ $2.7$ 2020
Top-Heavy IMF
MSs ($0.01M_{\odot}-150M_{\odot}$)
$0.02$ 11 $4.0\times 10^{6}$ $10$
MTG83F0.025k $1.7\times 10^{8}$ $4.1\times 10^{7}$ $2.3$ $4.1$ $24$ $0.025$ 11 $4.0\times 10^{6}$ $11$
In this section, we extend our studies by including three additional complexities in the simulation: components with a spectrum of masses, stellar evolution processes (see details in Section 2.4), and the additional mass growth of the MBH from accreting gas released by mass loss during stellar evolution. The initial conditions of the models investigated in this section are shown in Table 2.
We consider two types of IMFs. The first follows the Kroupa IMF (Kroupa, 2001) for MSs ranging from $0.01M_{\odot}$ to $150M_{\odot}$. The number distribution of stars in the Kroupa IMF follows $f(m_{\star})\propto m_{\star}^{\alpha}$, where $\alpha=-0.3$ if $m_{\star}\in(0.01,0.08)M_{\odot}$, $\alpha=-1.3$ if $m_{\star}\in(0.08,0.5)M_{\odot}$, and $\alpha=-2.3$ if $m_{\star}\in(0.5,150)M_{\odot}$. Observations in our Galactic Center suggest that the NSC may follow a top-heavy IMF with $\alpha=-1.6\sim-0.45$ (Bartko et al., 2010; Lu et al., 2013). Therefore, we implement a second IMF model that is similar to the Kroupa IMF but adopts $\alpha=-1.6$ for $m_{\star}\in(0.5,150)M_{\odot}$. For all models, we assume a metallicity of $Z=0.02$.
The diffusion coefficients are calculated by GNC in discrete mass bins (Paper II). For all models in Table 2, we define mass bin edges at $0.01M_{\odot}$, $0.1M_{\odot}$, $0.5M_{\odot}$, $1M_{\odot}$, $2M_{\odot}$, $4M_{\odot}$, $8M_{\odot}$, $12M_{\odot}$, $24M_{\odot}$, $32M_{\odot}$, $64M_{\odot}$, $128M_{\odot}$, and $150M_{\odot}$, resulting in $12$ mass bins for the evaluation. We find that this number of mass bins is sufficient for the convergence of the model results.
A significant amount of mass is expected to be lost during stellar evolution. For example, when a star evolves from the main sequence to the giant branch and finally to a compact object, a significant amount of its mass is lost during these transitions. We assume that a fraction $f_{\rm ma}$ of the lost mass is added to a gas reservoir, which can be subsequently accreted by the central MBH. Since the majority of the MBH’s mass growth is expected to come from stellar evolution mass loss (e.g., Freitag & Benz, 2002; Freitag et al., 2006), the factor $f_{\rm ma}$ plays a crucial role in determining the long-term evolution of the MBH mass and the structure of the NSC.
There are large uncertainties in $f_{\rm ma}$, as it is possible that some gas may fragment and form stars before falling into the center and being accreted by the MBH. Murphy et al. (1991) adopt a fraction of $100\%$. Freitag et al. (2006) make an ad hoc assumption of $f_{\rm ma}=6.53\%$ to reproduce $M_{\bullet}/M_{\rm cl}=0.05$ after $\sim 10$ Gyr. In this study, we explore $f_{\rm ma}$ ranging from 0 to $1$, as shown in Table 2.
We primarily focus on models that can reproduce Milky-Way-like NSCs with masses of $\sim 4\times 10^{7}M_{\odot}$ by $12$ Gyr. These models start with cluster masses of $M_{\rm cl}=7\times 10^{7}M_{\odot}$ or $1.7\times 10^{8}M_{\odot}$, depending on the IMF. We also study more massive NSCs with initial masses of $M_{\rm cl}=10^{9}M_{\odot}$ and small NSCs with $M_{\rm cl}=2\times 10^{6}M_{\odot}$.
We first provide detailed analyses of the co-evolution of the NSC and the MBH, including the evolution of stellar components (Section 3.2.1), the NSC structure (Section 3.2.2), TDEs (Section 3.2.3), and MBH mass growth (Section 3.2.4). These results are critical for understanding the formation and evolution of EMRIs, as together they establish the environmental background where EMRIs form. In Section 3.2.5, we focus specifically on models that can reproduce the Milky-Way NSC.
We then analyze the evolution of EMRI events in Section 3.2.6. The expected mass, eccentricity, and orbital inclination distributions in the LISA band will be discussed in Section 3.2.7 and 3.2.8, respectively.
3.2.1 The evolution of stellar populations
Figure 6: First panel: Evolution of the number of different stellar objects assuming a Kroupa IMF, taken from the results of Model MKG82F0.06k in Table 2. Second panel: Similar to the first panel but for the number ratio of different stellar objects to MSs. Third and fourth panels: Similar to the first and second panels but for a top-heavy IMF, taken from the results of Model MTK82F0.02k.
The numerical evolution of MSs, Post-MSs, BDs, and compact objects plays an important role in understanding the evolution of NSCs and EMRIs. The evolution of the stellar populations for the Kroupa IMF and the top-heavy IMF is shown in Figure 6.
We can see that for the Kroupa IMF, the number ratios of SBHs, NSs, WDs, BDs, and Post-MSs to MSs are approximately $1.3\times 10^{-3}$, $6.3\times 10^{-3}$, $0.11$, $0.91$, and $5.4\times 10^{-3}$ at $12$ Gyr, respectively. Since SBHs and NSs are remnants of massive MSs ($\gtrsim 8M_{\odot}$) with relatively short lifetimes, their populations rapidly peak around $\sim 0.01-0.1$ Gyr and then remain nearly constant. In contrast, the WD population grows continuously over cosmic time due to the evolution of low-mass stars. The Post-MS population generally increases until $\sim 1$ Gyr, after which it exhibits minor variations due to complex post-main-sequence evolution.
If we alternatively adopt a top-heavy IMF with $\alpha=-1.6$ for $m\gtrsim 0.5M_{\odot}$, the evolution is similar, but the populations of Post-MSs and compact objects increase significantly. By $12$ Gyr, the number ratios of SBHs, NSs, WDs, BDs, and Post-MSs to MSs become $0.041$, $0.059$, $0.29$, $0.85$, and $8.6\times 10^{-3}$, respectively.
3.2.2 The evolution of the NSC’s structure
Figure 7: Top left panel: Cluster mass $M_{\rm cl}$ and effective radius $r_{\rm eff}$ of some models in Table 2. For each model, the point on the left represents the initial value, and the point on the right represents the value at $12$ Gyr. The gray circles with error bars indicate observations of nearby galaxies from Georgiev et al. (2016). The cyan and dark-green solid straight lines represent the fitted correlation between $M_{\rm cl}$ and $r_{\rm eff}$ by Georgiev et al. (2016) for late- and early-type galaxies, respectively. The black empty star marks the value in our Galactic Center. Top Right panel: The evolution of the mean density of the cluster, $M_{\rm cl}/r_{\rm eff}^{3}$, as a function of the mass of the MBH, $M_{\bullet}$. Bottom left panel: The evolution of $r_{\rm eff}$ as a function of time. The shaded region indicates the current constraints on the Milky Way NSC’s effective radius ($4.2$ $r_{\rm eff}<7.2 pc$) (Neumayer et al., 2020). Bottom right panel: Density slope indices of different stellar objects versus the present-day mass of the MBH. Symbols for each model have the same color as the lines in the middle panel. The slope indices are calculated based on the densities between $10^{-3}r_{\rm eff}<r<0.1r_{\rm eff}$.
Figure 8: The mass growth of the MBH (first and third columns of panels) and the event rate evolution of EMRIs and TDEs (second and fourth columns of panels) for different models in Table 2. For each model, the legends are similar to those described in Figure 5 (additionally, ”TD(Post-MS)” means TDEs of Post-MSs). The event rates for each model have been smoothed via a moving average over every five or ten data points.
The evolution of the effective radius ($r_{\rm eff}$), the cluster’s stellar mass ($M_{\rm cl}$), and the mean density of the cluster ($M_{\rm cl}/r_{\rm eff}^{3}$) can be found in Figure 7. The MBH mass and the effective size of the cluster at $12$ Gyr can be found in Table 2.
We can see that NSCs exhibit significant size expansion and a decrease in mean density over cosmic time. The size evolution of an NSC is mainly affected by three factors:
  • Mass loss from the cluster due to stellar evolution, which weakens the stellar gravitational potential and causes stellar orbits to expand. As listed in Table 2, the cluster loses approximately $40\%$ of its mass over $12$ Gyr for a Kroupa IMF or $\sim 75\%$ for a top-heavy IMF, with most of the mass loss occurring within the first $100$ Myr. The larger the mass loss relative to the NSC’s mass, the more pronounced the expansion.
  • Stellar relaxation processes, which cause the cluster to expand (see Merritt (2009) and Section 4.2.1 of Paper II). The expansion also facilitates the escape of particles from the system. By $12$ Gyr, we find that the mass that has escaped from the NSC is $1\%$, $\sim 3\%$, and $\sim 15\%$ of the stellar mass for massive NSCs (e.g., model MKG91F0k), Milky-Way-like NSCs (e.g., model MKG83F0k), and small NSCs (e.g., model MKG71F0k), respectively.
  • When the $e$-folding timescale of MBH mass growth–the time required for its mass to increase by a factor of $\sim 2.7$, also known as the Salpeter timescale–is much shorter than the local relaxation timescale, the rapid growth of MBH’s mass can slow or even reverse the size expansion of the cluster. In this case, the conservation of radial action (see Section 2.3 in Paper II) causes stellar objects to sink toward the center, increasing the cluster’s density and reducing its effective size (this effect is more apparent for model MKG83F1k in the bottom left panel of Figure 7).
Thus, the typical evolution of an NSC’s size proceeds as follows: In early epochs (<100<100 Myr), stellar mass loss accelerates the size expansion. When the gas reservoir is large enough to sustain Eddington-limited accretion, the MBH’s mass growth is fast, slowing down or even reversing the expansion process. However, once the MBH mass is so large that the gas reservoir can no longer sustain Eddington-limited accretion, its growth rate declines, and stellar relaxation processes become dominant. This transition ultimately leads to cluster expansion.
At $12$ Gyr, the density slope of SBHs, averaged in the region $r/r_{\rm eff}\in(0.01,0.1)$, is $\sim-2.0$ (bottom right panel of Figure 7). In contrast, lighter objects, e.g., MSs, NSs, WDs, and BDs, exhibit a much shallower density profile, with slope indices ranging from 1.5-1.5 to 0.5-0.5. These results are found to be insensitive to the initial conditions of the cluster. Thus, the present-day density profile of a cluster with a spectrum of masses remains similar to the cases where the cluster contains only two components (stars and SBHs), as explored in Paper I.
3.2.3 The evolution of TDE rates
It is necessary to investigate the evolution of TDEs, as they can contribute a non-negligible fraction of mass to the cosmological mass growth of MBHs. Figure 8 shows the cosmological evolution of TDEs of MSs, BDs, and Post-MSs in different models (dotted lines in the panel). Similar to Paper II, we find that the TDE rates of both MSs and BDs evolve depending on the ratio between the gravitational influence radius of the MBH, $r_{h}$ (defined by $M_{\star}(<r_{h})=2M_{\bullet}$, where $M_{\star}(<r)$ is the enclosed stellar mass within $r$), and the cluster’s effective radius, $r_{\rm eff}$. If initially $r_{h}\lesssim 0.08r_{\rm eff}$, the rates evolve in three distinct phases: (1) an initially high rate that decreases with time; (2) a subsequent rise driven by the rapid, Eddington-limited mass growth of the MBH; and (3) a decline again once the MBH mass is sufficiently large that $r_{h}>0.1r_{\rm eff}$. In contrast, if initially $r_{h}\gtrsim 0.08r_{\rm eff}$, the evolution skips the first two phases and proceeds directly to phase 3.
The bottom right panel of Figure 11 shows the TDE rates of MSs averaged over the early universe (<1<1 Gyr) and later epochs ($1\sim 12$ Gyr). There is a systematic decrease in TDE rates over cosmic time.
For Post-MSs, they do not follow the above behavior, as their TDE rates also depend on their evolving population size (see Figure 6) and the evolving distribution of their stellar radii. When including these complexities, Figure 8 shows that their TDE rates generally exhibit an initial peak and then gradually decline over time.
The TDE rates of MSs are generally higher than those of BDs, which are in turn higher than those of Post-MSs. The TDE rates of BDs and Post-MSs are one or two orders of magnitude lower.
In all models, the peak TDE rates of MSs are usually $10^{-3}\sim 10^{-2}$ yr-1 for massive NSCs ($M_{\rm cl}=10^{9}M_{\odot}$) and $10^{-4}\sim 10^{-3}$ yr-1 for Milky-Way-like or smaller NSCs ($M_{\rm cl}=2\times 10^{6}M_{\odot}$). However, at $12$ Gyr, the rates reduce to $10^{-4}\sim 10^{-3}$ yr-1  for massive NSCs, $10^{-5}\sim 10^{-4}$ yr-1 for Milky-Way-like NSCs, and $10^{-7}\sim 10^{-6}$ yr-1  for smaller NSCs. These results are generally consistent with Paper II. However, while Paper II assumed equal-mass MSs of $1M_{\odot}$, we find here that the majority of MS-TDEs involve stars with masses of $\sim 0.5M_{\odot}$, which is two times smaller.
3.2.4 Mass growth of the MBH
The mass growth of the MBH over $12$ Gyr in different models can be found in Figure 8. In the early Universe, MBHs typically experience a phase of rapid mass growth through Eddington-limited accretion, lasting approximately 0.10.50.1-0.5 Gyr. This phase is mainly sustained by gas supplied from TDEs and stellar mass loss (if included). The duration of this phase depends on both the MBH’s gas consumption rate (which is higher for a more massive MBH) and the gas supply rate.
The mass increase of the MBH at $12$ Gyr is approximately the sum of $\delta M_{\bullet,\rm td}$ (mass contributed by accreting gas from TDEs) and $\delta M_{\bullet,\rm ml}$ (mass contributed by stellar mass loss). Then the final mass of the MBH at $12$ Gyr is
$M_{\bullet,f}$ $\simeq\delta M_{\bullet,\rm td}+\delta M_{\bullet,\rm ml}+M_{\bullet,i}$ (10)
$\simeq\langle m_{\star}\rangle_{\rm td}\bar{R}_{\rm td,MS}T_{12\rm\,Gyr}+M_{\rm cl,i}f_{\rm ml}f_{\rm ma}+M_{\bullet,i}$
where
  • $\langle m_{\star}\rangle_{\rm td}\sim 0.5M_{\odot}$ is the mean mass of TDEs (see Section 3.2.3);
  • $\bar{R}_{\rm td,MS}$ is the mean event rate of MS-TDEs;
  • $T_{12\rm\,Gyr}=12$ Gyr is the time of accretion;
  • $M_{\rm cl,i}$ is the initial mass of the cluster;
  • $f_{\rm ml}=0.4\sim 0.7$ (see Section 3.2.2) is the fraction of stellar mass lost relative to the initial cluster mass;
  • $f_{\rm ma}$ is the assumed fraction of mass loss that is added to the gas reservoir, which can be further accreted by the MBH.
  • $M_{\bullet,i}=10^{4}M_{\odot}$ is the initial MBH mass ($M_{\bullet,i}=10^{4}M_{\odot}$).
According to Section 3.2.3 and Figure 8, the mean MS-TDE rates are generally $\bar{R}_{\rm td,MS}\sim 10^{-3}$ yr-1, $\sim 10^{-4}$ yr-1, and $\sim 10^{-6}$ yr-1  for massive NSCs ($M_{\rm cl,i}=10^{9}M_{\odot}$), Milky-Way-like NSCs, and smaller NSCs ($M_{\rm cl,i}=2\times 10^{6}M_{\odot}$), respectively. The total mass contributed by TDEs over $12$ Gyr is typically $\delta M_{\bullet,\rm td}\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. These results are consistent with the final MBH mass for models with $f_{\rm ma}=0$, e.g., models MKG91F0k, MKG82F0k, and MKG71F0k.
Note that since $\langle m_{\star}\rangle_{\rm td}\sim 0.5M_{\odot}$, the mass contributed by TDEs is about two times smaller compared to the models of Paper II, which assume MSs of $1M_{\odot}$.
Similar to Paper II, we find that the smaller the initial effective radius ($r_{\rm eff,i}$) of the NSC, the larger the rate of MS-TDEs ($R_{\rm td,MS}$), and thus the larger the final MBH mass at $12$ Gyr.
The mass contributed by stellar mass loss, $\delta M_{\bullet,\rm ml}$, depends critically on the assumed value of $f_{\rm ma}$. From Table 2, we can see that for models assuming $f_{\rm ma}\geq 0.1$, the final MBH mass is dominated by $\delta M_{\bullet,\rm ml}$, and the MBH-to-NSC mass ratio can be up to $\sim 0.1$ ($\sim 0.8$) if we assume $f_{\rm ma}=0.1$ ($f_{\rm ma}=1$).
3.2.5 Reproducing the Milky Way’s NSC
Figure 9: The spatial distribution of stellar objects illustrated based on model MKG82F0.06k. Stellar objects are assumed to be spherically distributed. For clarity, the number of MSs, Post-MSs, and SBHs is suppressed by a factor of $10^{4}$, $500$, and $50$, respectively. The top left panel shows the initial distribution of MSs (all zero-age MSs; BDs are not plotted for clarity), while the rest of the panels show the distribution of MSs, Post-MSs, and SBHs at $12$ Gyr. In all panels, filled colored circles represent individual stellar objects, with their sizes proportional to the stellar radii and their color representing the stellar mass according to the color bar on the right of each panel (in units of solar masses). The largest MS and Post-MS have radii of $20R_{\odot}$ and $226R_{\odot}$, respectively.
Figure 10: Simulation results for different stellar objects at $12$ Gyr for model MKG82F0.06k. Left panel: The mass density distribution as a function of radius. Middle panel: The cumulative stellar mass distribution as a function of radius. The data with error bars represents the constraints observed in the Milky Way’s NSC, while the shaded region marks the observational constraints on the total mass of the Milky Way’s NSC ($2.1\sim 4.2\times 10^{7}M_{\odot}$) (Schödel et al., 2009, 2014; Feldmeier et al., 2014; Chatzopoulos et al., 2015; Fritz et al., 2016; Feldmeier-Krause et al., 2017; Magorrian, 2019; Gravity Collaboration et al., 2024). Right panel: The slope index of the density as a function of radius. The gray solid line shows a reference slope index of $\gamma=-1.3$.
It is of particular interest to investigate models that reproduce the Milky Way’s NSC, for which we have the best observational constraints. From Table 2, we can see that models MKG81F0k and MKG82F0.06k(s) can reproduce the current mass of the MBH in the Galactic Center, i.e., $4\times 10^{6}M_{\odot}$ (Gillessen et al., 2009, 2017), and have an effective size in the range of $4.2$ $\sim 7.2 pc$ (Neumayer et al., 2020).
As discussed in Section 3.2.4, the mass growth of the MBH is contributed mainly by both TDEs of MSs (which can be regulated by the initial effective size of the cluster, $r_{\rm eff,i}$) and mass loss from stellar evolution (which depends critically on the assumed value of $f_{\rm ma}$). Model MKG81F0k ignores stellar mass loss ($f_{\rm ma}=0$) and thus requires a compact initial cluster size ($r_{\rm eff,i}=1.6$ pc) to reproduce the observed MBH mass at $12$ Gyr solely via TDEs. Model MKG82F0.06k includes stellar mass loss with an assumed value of $f_{\rm ma}=0.06$ and thus can initially have a slightly larger cluster size.
Note that to reproduce a Milky-Way-like NSC, $f_{\rm ma}$ cannot be larger than $\sim 0.1$. Otherwise, either the MBH mass or the effective size of the cluster would be larger than the observed values. This indicates that most of the gas released by stellar mass loss may not be efficiently accreted by the MBH. One possibility is that the gas collapses and forms stars before approaching the vicinity of the MBH.
Figure 9 illustrates the projected spatial distribution (assuming a spherical distribution), physical size, and masses of stellar objects for Model MKG82F0.06k. The initial cluster consists of zero-age MSs with masses ranging from $0.01-150M_{\odot}$. After $12$ Gyr, the MSs remaining in the cluster are those with masses $\lesssim 1M_{\odot}$, as more massive stars have already evolved into post-MSs or compact objects. At $12$ Gyr, the post-MSs are in the mass range of $0.6\sim 1M_{\odot}$, with much larger sizes of $1.5-230R_{\odot}$. From Figure 9, it is apparent that SBHs are more concentrated in the inner regions than MSs, mainly due to mass segregation effects.
The present-day density profile for model MKG82F0.06k is shown in Figure 10. This model can approximately reproduce the observed enclosed mass distribution (middle panel). For lighter components (MSs, Post-MSs, NSs, WDs, and BDs), they all have slope indices of $\sim-1.3$ within $\sim 0.1$ pc. The SBH population has a steeper slope index of $1.5\sim-2.5$. These results are roughly consistent with model M2G82 in Paper II, although the latter adopts a simple two-component model and lacks stellar evolution.
3.2.6 The evolution of EMRI event rates
Figure 11: The event rates of EMRIs ($R_{\rm EMRI}$) or TDEs ($R_{\rm TDE}$) averaged over the early universe ($t<1$ Gyr, empty red circles) and later epochs ($1$ Gyr<t<12<t<12 Gyr, empty green triangles) for some models in Table 2. In the bottom middle and right panels, the results of Model MKG82F0.06s are similar to those of Model MKG82F0.06k and are therefore not shown for clarity.
The evolution of EMRI event rates over $12$ Gyr can be found in Figure 8. For a clearer comparison of rates versus MBH masses across different models and cosmic times ( averaged over early epochs, the first $1$ Gyr, and later epochs, between 1121-12 Gyr), see Figure 11.
According to these results, we find that the evolution of EMRI rates is mainly affected by the following four factors:
  • The size expansion or contraction of the cluster (see also Section 3.2.2). The expansion of the cluster’s size due to either substantial stellar mass loss or relaxation processes leads to a decrease in the cluster’s density. As $R_{\rm EMRI}\propto M_{\rm cl}/r_{\rm eff}^{3}$ approximately, the EMRI rates for all stellar objects are expected to decline. However, if the growth of the MBH mass is sufficiently fast such that the local relaxation time is longer than the $e$-folding time of the MBH’s mass ($\sim 4.5\times 10^{7}$ yr), the shrinking of stellar orbits can slow down or even reverse the size expansion. If the latter occurs, the EMRI event rates can increase. This effect is particularly apparent in model MKG83F1k, where all EMRI rates show a very substantial increase around $\sim 0.5$ Gyr.
  • The evolution of the stellar population (see also Section 3.2.1). Almost all SBHs and NSs form in the early universe, and afterward, the sizes of their populations remain almost constant over cosmic time. Thus, their rates are affected mainly by the dynamical evolution of the cluster and the MBH. However, the rate evolution of WD-EMRIs is slightly complicated by the growing number of their populations over cosmic time (See Figure 6).
  • The IMF of the cluster. The IMF affects the evolution of the number and mass distribution of stellar populations over cosmic time (see Section 3.2.1). Different IMFs lead to different amounts of stellar mass loss, which affects the size expansion of the cluster and the mass growth of the MBH. These complications ultimately lead to differences in the observed rates and mass distribution of EMRIs.
  • The spin of the MBH. Around a maximally spinning MBH, the EMRI event rates of compact objects are usually $3\sim 4$ times higher than those around a non-spinning MBH. For MSs and BDs, however, the enhancement is less significant, as their loss cone sizes cannot be smaller than their tidal radii.
  • For BDs and MSs, they become EMRIs only if $M_{\bullet}\gtrsim 10^{5}M_{\odot}$ (see Section 3.1.2). The ISO pericenters of BDs and MSs are outside their tidal radii only if $M_{\bullet}\gtrsim 10^{6}M_{\odot}$ (see Figure 1). Around smaller MBHs, they are more likely to be tidally disrupted when passing through the loss cone. Thus, starting from a seed MBH of $10^{4}M_{\odot}$, BD-EMRIs and MS-EMRIs appear at later epochs of the Universe. BD-EMRIs and MS-EMRIs form more easily around more massive MBHs, as their tidal radii become relatively smaller compared to the ISO. However, how their rates evolve with time is also affected by the factors listed above.
Combining all the above complexities, we summarize and discuss the detailed evolution of individual stellar populations as follows:
  • SBH-EMRI events first appear at $10$ Myr and peak in the early universe ($\lesssim 0.1$ Gyr) with rates of $10^{-7}$ yr-1 $\sim 10^{-5}$ yr-1. After that, they gradually decrease by a factor of 5105-10 by $\sim 12$ Gyr.
    The evolution of these rates over cosmic time provides a crucial context for interpreting analytical steady-state models focused on the present epoch. For instance, Amaro Seoane et al. (2024) analytically investigated the population of Early-EMRIs (E-EMRIs)—sources in the long-lived early inspiral phase. They derived merger rates for the Galactic Center ranging from $10^{-6}$ to $10^{-5}$ yr-1 for $10M_{\odot}$ SBHs, assuming the system is currently relaxed.
    Our simulations for Milky-Way-like NSCs (e.g., models MKG81F0k and MKG82F0.06k) show peak SBH-EMRI rates of $\sim 3\times 10^{-6}$ yr-1 in the early universe ($\lesssim 1$ Gyr), which aligns well with the lower end of these analytical estimates. The simulations show a subsequent decline to $\sim 10^{-7}$ yr-1 by 12 Gyr, driven by long-term effects such as cluster expansion and the depletion of SBHs.
    The key concept in Amaro Seoane et al. (2024) is the steady-state number of sources currently observable in the LISA band, calculated by multiplying the merger rate by the long residence time ($T_{\rm res}\approx 1.85\times 10^{5}$ years for a $10M_{\odot}$ SBH with SNR >10>10). While our simulations track the long-term cosmological evolution, the analytical steady-state calculation provides an estimate for the present snapshot in time, relevant for LISA observations in the near future. The GNC simulations confirm the underlying dynamical processes—two-body relaxation feeding the GW-driven inspiral (Figure 2). If the current state of the Galactic Center supports the analytically derived merger rates (which are within an order of magnitude of our peak simulated rates), the prediction of a significant population of E-EMRIs (10-25 sources) currently in the band (Amaro Seoane et al., 2024) remains a plausible estimate for present-day observations, as the astrophysical conditions are not expected to change significantly over the residence timescale ($10^{5}$ years).
  • NS- and WD-EMRIs first appear slightly later than SBH-EMRIs. Their event rates initially peak at $\sim 0.1-1$ Gyr and then gradually decline over cosmic time. The decrease in WD-EMRI rates is not as significant as that of NS-EMRIs, as the number of WDs grows with time. Typically, the peaks of NS-EMRI (WD-EMRI) event rates are in the range of $10^{-9}\sim 10^{-7}$ yr-1($10^{-8}\sim 10^{-7}$ yr-1). The event rates of both NS- and WD-EMRIs are about 232-3 orders of magnitude lower than those of SBH-EMRIs.
  • For MS- and BD-EMRIs, they only appear when the MBH mass is larger than $\sim 10^{5}M_{\odot}$. After that, most models exhibit an increase in rates with MBH mass. In general, BD-EMRI and MS-EMRI event rates increase from $10^{-11}\sim 10^{-9}$ yr-1 for MBHs with $\sim 10^{6}M_{\odot}$ to $10^{-8}\sim 10^{-7}$ yr-1  for MBHs with $\sim 10^{7}M_{\odot}$. In most cases, MS-EMRI event rates are slightly higher than those of BD-EMRIs, mainly because MSs are more abundant and have larger masses. Models MKG83F1k and MKG91F1k grow their MBH mass quickly to $10^{7}M_{\odot}$, and thus their rates peak in the early universe. When the phase of rapid mass growth ends, size expansion starts to dominate, and the rates decrease with time.
    BD-EMRIs has been specifically investigated analytically by Amaro-Seoane (2019), where they are termed X-MRIs due to the extreme mass ratio ($\sim 10^{8}$). Our simulations confirm the premise of that work: BDs around Milky-Way-like MBHs can indeed survive tidal disruption (as $r_{\rm td}<r_{\rm ISO}$, see Figure 1) and successfully form X-MRIs.
The rates of SBH-EMRIs are orders of magnitude larger than those of other stellar types, which can be understood from the analysis of the critical radius $a_{\rm crit}$ discussed in Section 3.1.2.
For the models that can reproduce the Milky Way’s NSC, i.e., models MKG81F0k and MKG82F0.06k, the EMRI rates of SBHs, NSs, WDs, BDs, and MSs are approximately $10^{-7}$ yr-1, $5\times 10^{-9}$ yr-1, $2\times 10^{-8}$ yr-1, $10^{-9}$ yr-1, and $4\times 10^{-9}$ yr-1 by $12$ Gyr, respectively.
Amaro-Seoane (2019) predicted a merger rate of $\dot{\Gamma}_{\rm X-MRI}\sim 10^{-6}\,\textrm{yr}^{-1}$ for the Galactic Center, assuming $\beta=1.5$ for BDs and $\gamma=1.75$ for SBHs. Our simulations for Milky-Way-like models (e.g., MKG82F0.06k) yield BD-EMRI rates of $\sim 10^{-9}$ yr-1 at 12 Gyr. This discrepancy arises for three reasons: (1) Amaro-Seoane (2019) assumes a steady solution and no MBH mass growth, so it is more appropriate to compare the results with the initial rates of model M5_2 in Section 3.1.3. As shown in the bottom panel of Figure 4, the initial BD-EMRI rate is around $\sim 10^{-7}$ yr-1, consistent with the results of Amaro-Seoane (2019) within one order of magnitude. (2) In GNC, the rates are time-dependent, and are mainly affected by the cluster expansion, the mass growth of MBH and the assumed IMF (see also ection 2.6). The expansion of the cluster reduces the central density of BDs, leading to lower rates of EMRIs. (3) By $12$ Gyr, the density slope of BDs in GNC is $\beta\sim 1.3$ (See right panel of Figure 10), which is flatter than the assumed slope of $1.5$. A flatter density profile results in smaller rates of EMRIs (See Equation 9).
However, the central argument of Amaro-Seoane (2019) concerns the steady-state population observable by LISA. Given the extremely long residence time of X-MRIs (up to $2\times 10^{6}$ years in the LISA band), they calculated a population of $\sim 20$ observable sources in the Galactic Center, assuming the system is currently relaxed. While the long-term evolution in our simulations leads to lower present-day merger rates than the idealized analytical model, the dynamical pathway for forming X-MRIs is validated. The analytical estimate of the steady-state population remains a relevant prediction for the current epoch—the timeframe when LISA will operate—provided the relaxation processes in the Galactic Center today are efficient enough to support the required influx, as the conditions will not change significantly over the residence timescale.
The model with a top-heavy IMF (MTG83F0.025k) has very substantial mass loss ($\sim 75\%$ of the total stellar mass), and thus the size expansion is quite significant (see bottom left panel of Figure 7). However, it also produces more SBHs and NSs over cosmic time compared to the case adopting a Kroupa IMF. Combining these two factors, the rates of SBH-EMRIs and NS-EMRIs are comparable to those in the Kroupa IMF case, while the rates of EMRIs involving other stellar objects are lower.
3.2.7 The mass and eccentricity distribution of EMRIs in the LISA band
Figure 12: The mass-eccentricity joint distribution of EMRI events in Models MKG82F0.06s, MKG82F0.06k, MTG83F0.025k, and MKG91F0k. The distribution is averaged over 55 $\leq t\leq 12 Gyr$. $\langle M_{\bullet}\rangle$ shows the mean mass of the MBH during this period. $e_{0.1\rm mHz}$ is the eccentricity of EMRIs when their orbital periods are $10^{4}$ s. The color contour represents the number of samples per unit of $m$ (or $\log m$) and $e_{0.1\rm mHz}$. Panels from top to bottom show the results for SBH-EMRIs, NS- and WD-EMRIs, and MS- and BD-EMRIs, respectively. The solid cyan lines represent $e_{\rm 0.1mHz,max}$ (Equation 12) for $r_{\rm ISO}(i=0^{\circ})=2.12r_{\rm g}$ if the model adopts $a=1$, and $r_{\rm ISO}=8r_{g}$ if it adopts $a=0$. In the bottom panels, the dashed cyan lines show $e_{\rm 0.1mHz,max}$ (Equation 12) based on the tidal radii of MSs or BDs.
Figure 13: The normalized eccentricity distribution of EMRIs for some models in Table 2. Lines in different colors are results for different stellar objects. The solid cyan vertical line in each panel shows the position of $e_{\rm 0.1mHz,max}$ (Equation 12) for each model.
Here we investigate the present-day (at $12$ Gyr in the simulation) mass and orbital eccentricity distributions of EMRIs in the LISA band. We focus on the eccentricity of EMRIs ($e_{\rm 0.1mHz}$) when their orbital frequencies have decayed to $f_{\rm orb}=0.1$ mHz ($P=10^{4}$ s). Their primary GW emissions are then around $0.2$ mHz, which is within LISA’s sensitivity range.
The characteristic orbital semi-major axis of these EMRIs is
$a_{\rm 0.1mHz}\sim 18.45~{\rm rg}\times\left(\frac{P}{10^{4}~{\rm s}}\right)^{1/3}\left(\frac{M_{\bullet}}{4\times 10^{6}~M_{\odot}}\right)^{-2/3}.$ (11)
The maximum value of the eccentricity ($e_{\rm 0.1mHz,max}$) is limited by the size of the loss cone:
$1-e_{\rm 0.1mHz,max}=0.434\times\frac{r_{\rm p,lc}}{8\,{\rm rg}}\left(\frac{M_{\bullet}}{4\times 10^{6}~M_{\odot}}\right)^{2/3},$ (12)
where $r_{\rm p,lc}=\max(r_{\rm td},r_{\rm ISO})$.
For a maximally spinning (non-spinning) Milky-Way-like MBH, the highest eccentricity for compact objects can be up to $\sim 0.88$ ($\sim 0.56$). For BDs and MSs, $e_{\rm 0.1mHz,max}$ is determined not only by the ISO but also by their tidal radii (see bottom panels of Figure 12). If the loss cone is determined by their tidal radii, $e_{\rm 0.1mHz,max}\sim 0.81$, which is independent of $M_{\bullet}$.
Note that for an MBH with $a=1$ ($a=0$) and $M_{\bullet}\gtrsim 10^{8}M_{\odot}$ ($\gtrsim 1.4\times 10^{7}M_{\odot}$), all EMRIs will have already fallen into the MBH before their orbits decay to $f_{\rm orb}=0.1$ mHz. Similarly, for stellar objects in retrograde orbits, they cannot become EMRIs at $0.1$ mHz if they are around maximally spinning MBHs with $M_{\bullet}\gtrsim 8\times 10^{6}M_{\odot}$.
Figure 12 and 13 show the mass-eccentricity joint distribution and the orbital eccentricity distribution of EMRIs for some models in Table 2, respectively.
For SBH-EMRIs, their masses are in the range of $5-23M_{\odot}$. The majority of them are around $8M_{\odot}$ if assuming a Kroupa IMF, or $\sim 20M_{\odot}$ if assuming a top-heavy IMF. For Milky-Way-like or smaller MBHs, SBH-EMRIs in the LISA band are more likely to be found with eccentricities near $e_{\rm 0.1mHz,max}$. However, for massive MBHs ($\gtrsim 10^{7}M_{\odot}$), the eccentricity distribution of SBH-EMRIs is almost flat.
The eccentricity distributions of NS- and WD-EMRIs are similar to those of SBH-EMRIs, although they are more concentrated towards the higher end. Unlike for SBH-EMRIs, this is true even for very massive MBHs. For all compact objects, the bias towards high eccentricities is most pronounced around small MBHs (e.g., $\sim 10^{5}M_{\odot}$, see the bottom right panel of Figure 13).
MSs can become EMRIs only if their masses are in the range of $0.1M_{\odot}\lesssim m_{\star}\lesssim 1M_{\odot}$. More massive MSs have larger physical sizes and are thus all tidally disrupted before becoming EMRIs. Their eccentricity distribution is more concentrated towards low eccentricities compared to other stellar objects.
BDs are more likely to become EMRIs if they are more massive. For less massive BDs (e.g., $m\lesssim 0.05M_{\odot}$), their physical sizes are larger, and their GW orbital decay is so weak that the critical distances $a_{\rm crit}$ defined in Section 3.1.2 are much smaller. Below the maximum eccentricity $e_{\rm 0.1mHz,max}$, their eccentricity distribution peaks at lower eccentricities compared to compact objects.
The relatively high eccentricities of EMRIs observed in the LISA band can be understood as follows. Most EMRI samples form in the regime bounded by $T_{\rm GW}(x,j)\lesssim 0.1T_{\rm rlx}(x,j)$ and $j_{\rm lc}<j<1$. Since stellar objects are more abundant in the outer parts of the cluster, within this bounded regime, EMRIs are more likely to have higher eccentricities. The details of the distribution are then regulated by the position of $a_{\rm crit}$, the spin of the MBH, and the radial distribution of the stellar objects.
3.2.8 The inclination distribution of EMRIs in the LISA band
Figure 14: Similar to Figure 13 but for the orbital inclination distribution of EMRIs.
Figure 14 shows the orbital inclination distribution of EMRIs. We can see that if $a=0$, the inclinations of all EMRIs are uniformly distributed, simply because in this case, the loss cone size is independent of the orbital inclination.
However, if $a=1$, EMRIs of compact objects are more likely to have lower inclinations (i.e., they are more likely to be in prograde orbits). The reasons are mainly twofold: (1) ISOs with near-prograde orientations are smaller, so GW radiation is stronger and stellar objects more easily become EMRIs; (2) If $M_{\bullet}\gtrsim 8\times 10^{6}M_{\odot}$, stellar objects in near-retrograde orbits will have already fallen into the MBH before reaching $f_{\rm orb}=0.1$ mHz.
For BD- and MS-EMRIs, it is possible that for a fraction of them, the loss cone is determined by the ISO, while for the rest, it is determined by their tidal radii. For the latter subset of samples, their inclination distribution should be isotropic. Thus, for BD- and MS-EMRIs, the concentration of inclinations toward lower angles is generally less pronounced. If $M_{\bullet}\gtrsim 10^{7}M_{\odot}$, the loss cone of most BDs is determined by the ISO (see the bottom right panel of Figure 12), and thus their inclination distribution is similar to that of compact objects.
4 Discussion
By comparing the model results from Section 3.1 and the results presented in Section 4 of Paper II with those in Section 3.2, we can see that ignoring stellar evolution leads to significant underestimates of the cluster’s size expansion, the loss of the NSC’s total mass, and the mass growth of the MBH. Thus, our results demonstrate that stellar evolution must be included to achieve realistic simulations of NSC dynamics and MBH evolution, which is crucial for accurate predictions of TDE and EMRI properties.
Although our results provide substantial new insights into the evolution of NSCs, TDEs, and EMRIs, several important limitations should be noted. First, our assumption of a single burst of star formation in NSCs is an oversimplification, as observations suggest an extended star formation history in both our Galactic Center (e.g., Bartko et al., 2010) and nearby galaxies (Neumayer et al., 2020). However, as shown by  Nogueras-Lara et al. (2020) and Schödel et al. (2020), it is possible that $80\%$ of the current mass of the Milky Way’s NSC formed between $11\sim 13$ Gyr ago, suggesting that our simplified scenario of single-burst formation $12$ Gyr ago may not deviate substantially from reality.
Nevertheless, it will be necessary in the future to incorporate star formation processes into our method, which will require detailed modeling of the cosmic evolution of the star formation history of galaxies, the conversion efficiency of gas to stars, and the feedback mechanisms of the MBH’s accretion activity.
Our study primarily focuses on the cosmological evolution of EMRI formation rates. This contrasts with analytical studies that often assume a steady-state distribution of stars to calculate the current population of EMRIs in the LISA band, such as the Early-EMRIs (E-EMRIs) studied by Amaro Seoane et al. (2024) and the brown dwarf X-MRIs studied by Amaro-Seoane (2019). These analytical works highlight that the long residence time of these sources in the LISA band (from $10^{5}$ years for SBHs to $10^{6}$ years for BDs) can lead to a significant observable population even if the merger rates are low.
When comparing our time-dependent numerical results with these steady-state analytical predictions, we observe differences in the merger rates at 12 Gyr. The simulations show that long-term effects, such as stellar evolution mass loss and relaxation processes, lead to cluster expansion and a reduction in central density, which tends to suppress EMRI formation rates at late epochs compared to idealized analytical models.
However, the analytical models provide valuable estimates for the currently observable population, which is the relevant timeframe for the upcoming LISA mission. Provided the nuclear star cluster is currently relaxed, the steady-state analysis offers a snapshot of the expected EMRI population. The GNC simulations confirm the underlying dynamical mechanisms assumed in the analytical work. The analytical estimates of the steady-state population (Amaro Seoane et al. (2024), Amaro-Seoane (2019)) remain valid estimates (within an order of magnitude) for the present epoch, as the astrophysical conditions are not expected to change significantly over the relatively short residence timescales of the EMRIs (e.g., the next $10^{5}$ years), even though the simulations demonstrate that these conditions evolve over a Hubble time.
Another limitation is our omission of the bulge and other mass components surrounding the NSCs. For example, the current outer edge of a Milky-Way-like NSC in our simulation is around $\sim 1$ kpc, within which the enclosed dark matter mass can be up to $10^{9}M_{\odot}$, assuming a Milky Way NFW profile (McMillan, 2017). Including these additional components may suppress the further size expansion of the cluster, change the dynamics and leading to increase of both the TDEs and EMRI event rates.
Additionally, we have ignored the natal kicks of NSs and SBHs during their formation. Panamarev et al. (2019) suggest that natal kicks could eject a significant fraction of NSs from an isolated NSC. However, many of them may remain bound to the system when considering the additional gravitational potential created by the mass of the bulge and dark matter. We defer these more complex scenarios to future studies.
5 Conclusion
Extreme mass-ratio inspirals (EMRIs) are important targets for next-generation space-borne gravitational wave (GW) telescopes. Here, we investigate the dynamics and properties of EMRIs shaped by the co-evolution of massive black holes (MBHs) and nuclear star clusters (NSCs). We use our previously developed and well-tested Monte Carlo method, GNC, which can solve the self-consistent dynamics of NSCs with a mass-growing central MBH.
In order to investigate EMRI events, GNC has now been updated to include GW orbital decay, the loss cone of a spinning MBH, and stellar evolution. We adopt a Kroupa or top-heavy initial mass function (IMF) in the mass range of $0.01-150M_{\odot}$. The mass of the MBH grows not only by directly swallowing objects that fall into the loss cone but also by accreting gaseous material released by tidal disruption events (TDEs) and mass loss from stellar evolution.
We perform simulations over $12$ Gyr, investigating the density and size evolution of NSCs along with the mass growth of MBHs, EMRIs of stellar-mass black holes (SBHs), neutron stars (NSs), white dwarfs (WDs), main-sequence stars (MSs), and brown dwarfs (BDs), as well as TDEs of MSs, BDs, and post-main-sequence stars (Post-MSs).
We observe a significant loss of the cluster’s mass due to stellar evolution ($\sim 40\%$ for a Kroupa IMF and $\sim 75\%$ for a top-heavy IMF). The size of the NSC expands due to mass loss and relaxation processes. However, if the MBH is in a phase of Eddington-limited accretion and if the local relaxation timescale exceeds the Salpeter timescale, the size expansion can be slowed or even reversed.
Over $12$ Gyr, the mass growth of the MBH contributed by TDEs is typically $\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive (with a stellar mass of $10^{9}M_{\odot}$), Milky-Way-like, and smaller NSCs (mass of $2\times 10^{6}M_{\odot}$), respectively. The majority of MS-TDEs have masses of $\sim 0.5M_{\odot}$. Meanwhile, the mass contributed from stellar mass loss depends critically on the assumed value of $f_{\rm ma}$, which is the fraction of stellar mass loss that can be eventually accreted by the MBH. Assuming $f_{\rm ma}\geq 0.1$, the final MBH mass is dominated by the contribution from stellar mass loss. The MBH-to-NSC mass ratio can be up to $\sim 0.1$ ($\sim 0.8$) if assuming $f_{\rm ma}=0.1$ ($f_{\rm ma}=1$).
The evolution of MS- and BD-TDEs exhibits similar decline-rise-and-fall behaviors discussed in our previous work (Zhang & Amaro Seoane, 2025). For Post-MSs, their TDE rates generally decline continuously over cosmic time.
Reproducing a Milky-Way-like NSC requires $f_{\rm ma}\lesssim 0.1$ and an initial effective size for the cluster a few times smaller than the present value. This indicates that in our Galactic center, most of the gas released by stellar mass loss is not accreted by the MBH.
We find that the evolution of EMRI event rates is mainly affected by the size expansion or contraction of the cluster, the evolution of the stellar population, the IMF of the cluster, and the spin of the MBH. Rapid mass growth of the MBH can slow or even reverse the decline of EMRI rates. For BD- and MS-EMRIs, they exist only if the MBH mass is much larger than $\sim 10^{5}M_{\odot}$; otherwise, they are more likely to be tidally disrupted.
Typically, the rates of SBH-, NS-, and WD-EMRIs peak at early epochs ($\lesssim 1$ Gyr) at $10^{-7}-10^{-5}$ yr-1, $10^{-9}-10^{-7}$ yr-1, and $10^{-8}-10^{-7}$ yr-1, respectively, and then gradually decline by a factor of 2102-10 over cosmic time. As the populations of WDs increase continuously with time, their EMRI rates decline more slowly than those of SBH- and NS-EMRIs. MS- and BD-EMRI event rates typically increase with the mass of the MBH. By $12$ Gyr, their rates can be up to $10^{-8}\sim 10^{-7}$ yr-1  if the MBH mass can grow up to $\sim 10^{7}M_{\odot}$.
For the models that can reproduce the Milky Way’s NSC, the EMRI rates of SBHs, NSs, WDs, BDs, and MSs are approximately $10^{-7}$ yr-1, $5\times 10^{-9}$ yr-1, $2\times 10^{-8}$ yr-1, $10^{-9}$ yr-1, and $4\times 10^{-9}$ yr-1at $12$ Gyr, respectively.
We investigate the mass, eccentricity, and inclination distributions of EMRIs appearing in the LISA band (when the orbital period reduces to $10^{4}$ s). The masses of most SBH-, BD-, and MS-EMRIs are in the ranges of $5\sim 23M_{\odot}$, $0.05\sim 0.1M_{\odot}$, and $0.1\sim 1M_{\odot}$, respectively.
For compact objects, they tend to have high eccentricities in the LISA band, especially around Milky-Way-like or smaller MBHs ($\lesssim 4\times 10^{6}M_{\odot}$). If the MBH is maximally spinning, EMRIs are biased towards low orbital inclination angles (prograde orbits). In these two cases, the eccentricity and inclination distributions of both MS- and BD-EMRIs are usually distinct from those of compact objects.
Our study provides numerous new details about the evolution of NSCs, the mass growth of MBHs, and EMRIs, which will be useful for the analysis and understanding of EMRIs observed by LISA or other space-based GW observatories. However, our conclusions are still limited by some oversimplifications in our simulations, e.g., the omission of an extended star formation history in the cluster and the bulge and dark matter components surrounding the NSC. In the future, we will continue to incorporate more necessary recipes in our study to achieve a more comprehensive and realistic understanding of NSCs over cosmic time.
6 Acknowledgments
We thank Matteo Sadun Bordoni and Sebastiano Fellenberg for helpful discussions of this work. This work was supported in part by the National Natural Science Foundation of China under grant Nos. 12273006 and 11833007. This work was also supported in part by the Key Project of the National Natural Science Foundation of China under grant No. 12133004. The simulations in this work were performed partly at the TianHe-II National Supercomputer Center in Guangzhou.
Appendix A Solving for the Relativistic ISO
Figure A: Left panel: The pericenter distance $r_{p}$ (in units of $r_{g}$) of an innermost stable orbit (ISO) in Boyer-Lindquist (BL) coordinates, plotted as a function of the orbital semi-major axis, defined as $a_{\rm GR}=(r_{p}+r_{a})/2$, where $r_{a}$ is the apocenter distance. Both $r_{p}$ and $r_{a}$ are solved using the method described in Appendix A. The colored lines show the results for the case of $a=1$ with various inclination angles ii. The solid black line shows the result for a Schwarzschild MBH. Middle panel: The mapped Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian semi-major axis $a_{\rm NW}$, both defined by Equation B1. Right panel: The Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian inclination angle $i_{\rm NW}$, calculated for orbits with a semi-major axis of $a_{\rm NW}=10^{4}r_{g}$.
Consider a particle of mass $m$ orbiting a Kerr MBH of mass $M_{\bullet}$. Its spacetime position is expressed in Boyer-Lindquist (BL) coordinates $(r,\theta,\phi,t)$. Let $r_{a}$ and $r_{p}$ denote the apocenter (maximum) and pericenter (minimum) radial distances to the MBH, respectively. They are two of the roots of the following equation (in units where $M_{\bullet}=G=c=1$(Bardeen et al., 1972):
$R(r)=[(r^{2}+a^{2})-a\lambda]^{2}-(r^{2}-2r+a^{2})[\xi^{2}r^{2}+(\lambda-a)^{2}+q^{2}]=0$ (A1)
where $\xi=m/E$, $\lambda=L_{z}/E$, and $q^{2}=Q/E^{2}$. Here, $E$ is the energy, $L_{z}$ is the angular momentum component along the spin axis, and $Q$ is the Carter constant of the particle. $a$ is the dimensionless spin parameter of the MBH. For particles with $m\neq 0$, the above equation has four real roots, $r_{a}\geq r_{p}\geq r_{3}>r_{4}$. For an ISO, the condition $r_{p}=r_{3}$ must be satisfied, which implies:
$\left.\frac{dR(r)}{dr}\right|_{r=r_{p}}=0.$ (A2)
The turning points in $\theta$ are given by the roots of the following equation:
$\Theta(\mu)=q^{2}-\mu^{2}\left[a^{2}(\xi^{2}-1)+\frac{\lambda^{2}}{1-\mu^{2}}\right]=0,$ (A3)
where $\mu=\cos\theta$. We define
$\cos i=\frac{L_{z}}{\sqrt{L_{z}^{2}+Q}}=\frac{\lambda}{\sqrt{\lambda^{2}+q^{2}}}$ (A4)
as the ”inclination” of the orbit. One of the turning points in Equation A3 then satisfies:
$\mu_{+}^{2}=\cos^{2}(\pi/2-i)$ (A5)
The value of $\lambda$ is given by the solution to the following equation (Sopuerta & Yunes, 2011):
$\mathscr{C},\mathscr{B}]\lambda^{2}+[\mathscr{D},\mathscr{B}]\lambda+[\mathscr{E},\mathscr{B}]=0$ (A6)
where $x,y]=x(r_{a})y(r_{p})-y(r_{a})x(r_{p})$ and
$\mathscr{B}$ $=-(r^{2}-2r+a^{2})(r^{2}+\mu^{2}_{+}a^{2})$ (A7)
$\mathscr{C}$ $=-\frac{a^{2}\mu_{+}^{2}+r^{2}-2r}{1-\mu_{+}^{2}}$
$\mathscr{D}$ $=-4ra$
$\mathscr{E}$ $=(r^{2}+a^{2})(r^{2}+\mu_{+}^{2}a^{2})+2ra^{2}(1-\mu_{+}^{2})$
$\xi$ can be solved for using:
$\mathscr{B}\xi^{2}+\mathscr{C}\lambda^{2}+\mathscr{D}\lambda+\mathscr{E}=0$ (A8)
$q^{2}$ is then given by:
$q^{2}=\mu_{+}^{2}\left[a^{2}(\xi^{2}-1)\right]$ (A9)
So far, for given values of $r_{a}$, $r_{p}$, and $i$, we can solve for $\xi$, $\lambda$, and $q^{2}$ using Equations A6,  A8, and A9.
The remaining two roots of Equation A1 are given by
$r_{3}=\alpha+\sqrt{\alpha^{2}-\beta},~r_{4}=\alpha-\sqrt{\alpha^{2}-\beta}$ (A10)
where
$\alpha=-\frac{\xi^{2}}{1-\xi^{2}}-\frac{r_{a}+r_{p}}{2},~\beta=-\frac{a^{2}q^{2}}{1-\xi^{2}}\frac{1}{r_{a}r_{p}}$ (A11)
In general, the pericenter $r_{p}$ of an ISO for given values of eccentricity $e$, inclination $i$, and MBH spin $a$ can be solved numerically by the following steps:
  1. 1.
    Set the initial value of $r_{p}$ to $r_{p}=r_{\rm hz}=1+\sqrt{1-a^{2}}$.
  2. 2.
    Calculate the apocenter distance $r_{a}=r_{p}/(1-e)$. Then, solve for $\lambda$ using Equation A6.
  3. 3.
    Solve for $\xi$ and $q^{2}$ using Equations A8 and A9, respectively. Then, calculate $r_{3}$ and $r_{4}$ using Equation A10.
  4. 4.
    If $r_{p}$ is not equal to $r_{3}$, update $r_{p}$ using the rule $r_{p}\rightarrow{\rm max}[(r_{p}+r_{3})/2,r_{\rm hz}]$ and return to step 2.
The above steps are repeated until $r_{p}$ is sufficiently close to $r_{3}$. The resulting value of $r_{p}$ is the pericenter of the ISO.
For the case $a=0$ (a Schwarzschild black hole), $r_{p}$ and $\lambda$ can be expressed analytically as:
$r_{p}=\frac{2(3+e)}{1+e}r_{g},~\lambda={\rm sgn}(L_{z})\frac{(3+e)^{3/2}}{2^{1/2}(1+e)^{1/2}}$ (A12)
Thus, in the limit $e\rightarrow 1$, we have $r_{p}\rightarrow 4r_{g}$ and $|\lambda|\rightarrow 4$. For a circular orbit ($e=0$), $r_{p}=6r_{g}$. The left panel of Figure A shows the solution for $r_{p}$ for orbits around a Schwarzschild ($a=0$) or a maximally spinning ($a=1$) MBH.
Appendix B Mapping to a Newtonian Loss Cone
The loss cone derived in the previous section is formulated in BL coordinates within a relativistic framework. It cannot be directly applied to GNC, which is defined in a purely Newtonian context. Consequently, a mapping between the relativistic and Newtonian loss cones must be established.
This mapping is not unique, however, as there is freedom in the choice of relativistic coordinates. For example, the transformation can be performed using the original BL coordinates, the local non-rotating rest frame, or harmonic coordinates. Nevertheless, all these mappings should converge at sufficiently large distances from the MBH. As the majority of particles in GNC have apocenter distances $\gg r_{g}$, we perform the mapping at the orbital apocenter under the Schwarzschild metric. In this case, the Newtonian semi-major axis ($a_{\rm NW}$, in units of $r_{g}$), eccentricity ($e_{\rm NW}$), and inclination ($i_{\rm NW}$) of the orbit are given by
$a_{\rm NW}$ $=-\frac{1}{2E_{\rm NW}},$ (B1)
$e_{\rm NW}$ $=1-v^{2}r,$
$\cos i_{\rm NW}$ $=\frac{v_{\phi}}{v}=i,$
where
$E_{\rm NW}$ $=\frac{1}{2}v^{2}-\frac{1}{r},$ (B2)
$v$ $=(v_{\theta}^{2}+v_{\phi}^{2})^{1/2},$
$v_{\theta}$ $=\pm\frac{q}{r}\left(1-\frac{2}{r}\right)^{1/2}$
$v_{\phi}$ $=\frac{\lambda}{r}\left(1-\frac{2}{r}\right)^{1/2}$
Given the values of $\lambda$ and $q$ for an ISO from the previous section, along with an inclination angle $i$, we can obtain the Newtonian pericenter $r_{\rm p,NW}=a_{\rm NW}(1-e_{\rm NW})$ using Equation B1. For nearly parabolic orbits where $e_{\rm NW}\simeq e\rightarrow 1$, the Newtonian angular momentum approaches $L_{\rm NW}=(r_{\rm p,NW}(1+e_{\rm NW}))^{1/2}\rightarrow(2r_{\rm p,NW})^{1/2}$, which in turn approaches $\lambda$. For a Schwarzschild MBH, as shown at the end of Section A, $\lambda\rightarrow 4$ when $e\rightarrow 1$. Thus, we immediately find that $r_{\rm p,NW}=8r_{g}$. This is the commonly adopted value for the pericenter of an ISO around a Schwarzschild MBH in studies using Newtonian approximations. The solutions for $r_{\rm p,NW}$ for different MBH spin parameters $a$ and orbital inclination angles are shown in Figure A.